### v2

# EvoXplain: When Machine Learning Models Agree on Predictions but Disagree on Why Measuring Mechanistic Multiplicity Across Training Runs

Abstract
Machine learning models are primarily judged by their predictive performance, especially in applied settings. Once a model reaches a high accuracy, its explanation is often assumed to be both correct and trustworthy. However, this assumption raises an often overlooked question: when two models achieve high accuracy, do they depend on the same internal logic, or do they reach the same outcome via different—and potentially competing—mechanisms?
We introduce EvoXplain, a simple diagnostic framework that measures the stability of model explanations across repeated training. Rather than analysing the explanation of a single trained model, EvoXplain treats explanations as samples drawn from the stochastic optimisation process itself—without aggregating predictions or constructing ensembles—and examines whether these samples form a single coherent explanation or instead separate into multiple, distinct explanatory modes.
We evaluate EvoXplain on the Breast Cancer and COMPAS datasets across two widely used model classes: Logistic Regression and tree-based models (Random Forests). Although all models achieve high predictive accuracy, their explanations frequently exhibit clear multimodality. Even models commonly assumed to be stable—such as Logistic Regression—can produce multiple, well-separated explanatory basins under repeated training on the same data split. Importantly, these differences are not explained by hyperparameter variation or simple performance trade-offs.
EvoXplain does not attempt to select a “correct” explanation. Instead, it makes explanatory instability visible and quantifiable, revealing when averaged or single-instance explanations obscure the existence of multiple underlying mechanisms. This provides a practical check on interpretability claims and highlights a failure mode of common explanation-averaging practices used in ensemble and AutoML pipelines. More broadly, EvoXplain reframes interpretability as a property of a model class under repeated instantiation, rather than of any single trained model.
1 Introduction
Machine learning models are increasingly deployed in domains where explanations matter, including healthcare, public policy, and scientific discovery. Despite the fact that practitioners routinely retrain models during development and performance tuning, interpretability is typically assessed using a single trained instance—or by averaging explanations across runs—rather than by explicitly examining whether explanations themselves remain consistent under retraining. This practice implicitly assumes that high predictive accuracy corresponds to a single, stable way of reasoning (Lipton (2018); Rudin (2019)).
There have been sustained challenges to this assumption in the literature. Perhaps the most well-known is the Rashomon effect, which observes that distinct models may achieve similar predictive performance while relying on different internal logics (Breiman (2001); Marx et al. (2020)). Related work on underspecification highlights that modern learning pipelines often admit many solutions that are indistinguishable in terms of accuracy, leaving optimisation procedures to select arbitrarily among them (D’Amour et al. (2022)). Other studies have shown that explanations can vary substantially when models are retrained, or when features are correlated or perturbed, including documented failures of attribution methods under retraining, feature overlap, or input transformations (Kindermans et al. (2017); Nogueira et al. (2018); Kumar et al. (2020); Frye et al. (2020)). Collectively, these results suggest that explanation stability cannot be assumed a priori. However, existing work does not provide a practical way to quantify how much explanatory variability exists in a given setting, nor whether such variability is structured or merely diffuse.
To address this gap, we propose EvoXplain, a simple and easy-to-apply framework for assessing the stability of model explanations across repeated training. Rather than evaluating a single attribution vector, EvoXplain treats explanations as outcomes of stochastic optimisation and analyses them as a distribution induced by the training process itself. By clustering explanation samples obtained from repeated instantiations of the same model class—without aggregating predictions or constructing ensembles—EvoXplain determines whether training consistently yields a single explanatory mode or instead produces multiple, distinct explanatory regimes.
EvoXplain does not aim to decide which explanation is correct. Instead, it provides evidence about whether a model’s explanations are reliable or whether they depend sensitively on the particular optimisation trajectory taken during training. This perspective aligns with a growing view that interpretability should be evaluated empirically rather than assumed, a position increasingly reflected in discussions of responsible AI, model auditing, and governance (Arrieta and others (2020); Selbst and Barocas (2018); European Commission (2023)). By making explanation stability measurable, EvoXplain enables practical questions such as: does this model tell a consistent story, or does its reasoning change across retraining?
In this paper, we introduce the EvoXplain framework, demonstrate its application on high-stakes datasets, and discuss why explanation stability matters for scientific inference, auditing, and model governance. Our goal is to support interpretability practice by providing researchers and practitioners with a diagnostic tool for detecting when explanations are stable—and when multiple explanatory stories coexist despite similar predictive performance.
2 Related Work
2.1 Model Performance and Interpretability
A large proportion of studies in machine learning assess models primarily through the lens of accuracy. In the real world, interpretability is often categorized based upon the model being used. Linear models and decision trees, for instance, are often tagged as “interpretable” because practitioners are able to read and comprehend the parameters or structures involved. However, some recent contributions to the field argue that readability cannot be equated with the models achieving high predictive accuracy (Lipton (2018)). EvoXplain follows the same direction by considering interpretability as something that can be measured and verified rather than assumed.
2.2 Rashomon Effects and Predictive Multiplicity
Our work relates to the Rashomon effect in that different models can arrive at the same answer and yet rely on totally different paths or reasoning (Breiman (2001); Semengue et al. (2019)). Some recent research considers predictive multiplicity under settings of fairness and of unconstrained risk (Umbach et al. (2023); Marx et al. (2020); Fisher et al. (2019)), showing that the same risk scores can be justified through different models (Umbach et al. (2023)). Unlike these approaches, EvoXplain studies multiplicity within the same model class, trained on the same data, to see whether retraining leads to different explanation patterns.
2.3 Underspecification and Training Variability
Research on underspecification D’Amour et al. (2022) shows that training pipelines may converge to many possible solutions that perform equally well. EvoXplain complements this work by looking specifically at explanation behaviour. Instead of only asking whether predictions differ, EvoXplain checks whether internal reasoning differs even when predictions stay the same.
2.4 Feature Importance Instability
Previous studies indicate that SHAP, LIME, and gradients as explanation methods can produce instability due to retraining, dataset perturbations, or correlated features (Kindermans et al. (2017); Nogueira et al. (2018); Kumar et al. (2020); Agarwal et al. (2021)). Most of these studies focus on instability in some specific values. EvoXplain extends local noise analysis by detecting whether instability forms structured groups of explanations and how often each group appears.
2.5 Rashomon-Set Explanations in AutoML
A related line of work incorporates model multiplicity into post-hoc explanations by operating over Rashomon sets—collections of near-optimal models within a tolerance of the best model’s loss. Cavus et al. (2025) propose the Rashomon Partial Dependence Profile (Rashomon PDP), which aggregates partial dependence profiles across the Rashomon set and reports uncertainty bands to highlight regions of disagreement. While such approaches usefully acknowledge that multiple near-optimal models exist, they typically treat multiplicity as uncertainty around a single aggregated explanation. EvoXplain studies a different phenomenon: structured, discrete explanatory modes (basins) among equally accurate models, and shows that explanation aggregation can yield a ghost explanation that may not correspond to any realised model instance. In other words, rather than summarising variability around one explanation, EvoXplain detects and quantifies distinct mechanistic regimes within the same training pipeline.
2.6 Mechanistic Interpretability in Deep Models
The understanding of circuits and the internal functioning of large neural networks remains an ongoing area of research (Olah et al. (2020)). Specific techniques tend to rely on handcrafted analyses or other tools engineered for individual designs. On the other hand, EvoXplain focuses on using generic attribution outputs to determine mechanism-level differences.
2.7 Explainable AI for Science and Metafeature-Space Diagnostics
Recent work in explainable AI for scientific workflows studies where a fixed model tends to improve or degrade, by searching for subgroups of the feature space associated with changes in predictive behaviour. Soares et al. introduce Meta Subspace Analysis (MSA), which discovers exceptional subspaces and relates their metafeatures (e.g., distributional and dependency summaries) to model (mis)behavior via association rules (Soares et al. (2025)). This perspective is complementary to EvoXplain. MSA characterises heterogeneity across data subspaces for a given trained model, whereas EvoXplain characterises heterogeneity across trained model instances under the same pipeline. Thus, MSA does not overlap with EvoXplain’s core object of study (explanation basins over retraining), but motivates future combinations, such as measuring explanation multiplicity within subspaces identified as behaviourally exceptional.
2.8 Explanation Stability in Safety-Critical Machine Learning
Beyond interpretability research, the safety assurance community has emphasised the need for explicit evidence about model robustness under retraining. Habli and colleagues propose structured assurance methodologies for machine learning components, such as AMLAS and the “BIG Argument” for AI safety (Habli et al. (2025); Hawkins et al. (2021)), which require that system behaviour be justified and verified beyond predictive accuracy alone. EvoXplain contributes to this assurance perspective by offering a concrete diagnostic for explanation stability. When explanations diverge across retrainings of equally accurate models, this reflects a form of epistemic uncertainty that must be accounted for in any rigorous ML safety case. In this sense, EvoXplain does not replace assurance frameworks, but provides a measurable source of evidence that can be incorporated within them.
2.9 Summary
There is an acknowledgment of the following from the previous literature:
1.
Different distinguishable models are able to achieve the same level of performance.
2.
Optimization may lead the model to have multiple valid solutions.
3.
Explanations may differ during retraining.
What is absent is a simple way to quantify explanation variability, when the latter is present. More specifically, we need to understand explanation variability, its diversity, and its frequency. EvoXplain addresses this limitation by turning explanation stability to a visible property instead of an assumption.
3 Problem Formulation
We use interpretability methods to understand how models make decisions, but most analyses implicitly assume that a model has a single, stable explanation once trained. This assumption is not always justified. Due to randomness in the training process, model inductive biases, and model flexibility, it is possible for multiple different explanations to achieve similar predictive performance. A first step toward trustworthy interpretability is therefore to verify whether explanatory multiplicity exists at all.
Formally, consider a model class trained on a fixed dataset . Standard interpretability workflows examine a single trained instance and its corresponding attribution vector . In contrast, this work asks the following question: when is trained multiple times with independent random initializations, do the resulting attribution vectors converge to a single explanatory pattern, or do they separate into multiple distinct patterns?
Let denote independent training runs on , and let denote the corresponding explanation vectors. EvoXplain analyses the empirical distribution of to determine:
1.
whether explanations concentrate around a single dominant mode,
2.
whether they disperse into multiple stable explanatory modes, and
3.
how much support each mode receives.
In practice, we repeat this analysis across multiple train–test splits of to assess the robustness of explanatory behaviour under data resampling. However, the core object of interest remains the distribution of explanations induced by repeated training within a fixed data realisation.
EvoXplain offers an empirical method for answering these questions. Rather than assuming explanations are stable, it treats them as random samples generated by the training process and assesses their structure using clustering and entropy. This reframes interpretability as an empirical question: whether a model class tends to produce a single coherent explanation or multiple competing explanatory mechanisms under repeated instantiation.
Contributions
This work makes three main contributions.
First, we consider that explanation stability is something to be quantified. Instead of assuming that a model has a single reasoning way, we think of explanations as samples that can be analysed over several training runs.
Second, we introduce EvoXplain, a simple diagnostic framework that determines if explanations cluster into a single pattern or if they are dispersed across multiple clusters of different patterns. It also measures the occurrence frequency of each distinct explanatory pattern.
Third we apply EvoXplain across well-known and studied datasets using numerous widely used model classes. We observe that internal reasoning is not consistent even at high performance accuracy and this shows the need for explanation stability as a measure of model evaluation and supports interpretability approaches that examine the variability of explanatory stories a model class can produce.
4 Methods
EvoXplain treats explanations as outcomes that may vary across training runs, rather than as fixed properties of a single trained model or an average across multiple trained instances. Instead of assuming that a model class converges to a single mode of reasoning, EvoXplain empirically evaluates whether explanations are stable, whether multiple distinct explanatory patterns emerge, and how frequently each pattern occurs. The method follows a simple workflow: repeatedly train a model under fixed conditions, extract explanations, represent them in a common space, and analyse their stability.
4.1 Repeated Training and Attribution Sampling
We initially conducted the EvoXplain analysis using Gini-based feature importance and observed the same qualitative phenomenon of structured explanatory multiplicity across repeated training runs; we focus on SHAP in this paper because it is the attribution method most commonly used in deployed models and high-stakes decision-making contexts.
Let denote a model class (e.g., Logistic Regression or Random Forests). For each run , we train from an independent random seed while keeping the data split, preprocessing, and hyperparameters fixed. After each training run, we compute feature-level attributions using a chosen interpretability method (here, SHAP), yielding an attribution vector over the input features.
EvoXplain collects these attribution vectors,
and treats them as samples drawn from the explanation distribution induced by stochastic training. Importantly, EvoXplain does not aggregate predictions or construct ensembles; it analyses explanation variability arising solely from repeated instantiation of the same model class.
4.2 Representing Explanation Space
To enable comparison across runs, attribution vectors are normalised to remove scale effects and focus on relative feature importance. Specifically, we centre each vector using the mean attribution across runs and apply normalisation:
where is the mean attribution across runs. The resulting matrix represents the set of explanations as points in a shared geometric space.
4.3 Clustering and Basin Structure
To assess whether explanations converge to a single pattern or separate into multiple patterns, we apply -means clustering (chosen for simplicity and interpretability) for values of from to . For each , we compute the silhouette score and select
If , explanations concentrate around a single explanatory mode. If , EvoXplain identifies multiple explanatory basins corresponding to distinct explanation patterns reached across training runs.
4.4 Mechanistic Entropy and Basin Support
Let denote the cluster assignment of run . The empirical support of basin is defined as
We define mechanistic entropy as the Shannon entropy of this distribution,
normalised by (with if ), so that . Lower values indicate concentration around a dominant basin, while values near one indicate that training runs are broadly distributed across multiple explanatory basins.
4.5 Centroid Profiles and Mode Interpretation
Each explanatory basin is associated with a centroid
which represents the characteristic explanation pattern of that basin. Comparing centroid profiles reveals how and where the model’s reasoning diverges across runs, for example by emphasising different features or feature combinations.
4.6 Mechanistic Diagnosis
EvoXplain provides three core diagnostics:
1.
Mode count: whether explanations converge () or split into multiple modes.
2.
Mechanistic entropy: how training runs are distributed across explanatory modes.
3.
Basin profiles: the characteristic feature importance patterns of each mode.
These diagnostics do not identify a “correct” explanation; interpretation is left to domain experts and regulatory context.
4.7 Algorithm Summary
Clustering in explanation space is used here as a diagnostic tool rather than as a claim about the true number of underlying mechanisms. We adopt k-means with silhouette-based selection for simplicity and transparency, and because the resulting clusters are visually apparent in low-dimensional projections and consistent across independent train–test splits. Importantly, the presence or absence of explanation multiplicity does not hinge on fine-grained cluster boundaries: when training converges to a single explanatory regime (e.g., under fixed regularisation), explanations collapse to a single dense region and silhouette analysis consistently selects k = 1, whereas when multiple explanatory regimes are admissible, well-separated modes emerge with stable support across runs and splits. EvoXplain therefore does not rely on clustering to create structure, but uses clustering to quantify structure that is already present in the explanation distribution.
4.8 Theoretical Interpretation
EvoXplain can be understood through optimisation non-identifiability. Let denote the parameter space of a model class and its loss surface. In high-dimensional or weakly constrained settings, training may converge to multiple parameter configurations with near-optimal loss,
which we refer to as a mechanism manifold. Different training runs may select different points in and yield distinct attribution patterns even when predictive performance is similar.
Let map model parameters to attribution vectors. EvoXplain studies the empirical distribution
where clustering reveals groups of trained model instances with similar explanatory behaviour. Mechanistic entropy quantifies how training disperses across these groups. When , training concentrates near a single attractor; when , training repeatedly reaches different regions of the explanation manifold.
EvoXplain does not claim these basins correspond to true causal mechanisms. Rather, it demonstrates that optimisation and model capacity alone can induce multiple, stable explanatory behaviours.
4.9 Computational Considerations
For training runs and clustering over candidate values of , the computational cost scales as , where is the training time for one model and the number of features. Since training and attribution computations are independent across runs, EvoXplain is naturally parallelisable. In practice, a few hundred runs are often sufficient to reveal stable explanatory structure, while larger values of improve resolution.
4.10 Attribution Method Considerations
EvoXplain is agnostic to the choice of attribution method. Any technique producing feature-level relevance scores—such as SHAP, integrated gradients, gradientinput, permutation importance, or circuit-based attributions—may be used. While the geometry of explanatory basins may depend on the attribution method, the presence of instability often persists. EvoXplain therefore complements attribution tools by assessing the stability of their outputs rather than selecting among them.
4.11 Diagnostic Applications
EvoXplain can serve as a screening tool for interpretability audits, fairness assessments, scientific analysis, and model governance. By identifying cases where explanations fail to stabilise, it reveals when explanation use or aggregation should be treated with caution. In critical domains such as healthcare, criminal risk assessment, and public policy, EvoXplain can highlight when additional scrutiny or domain-specific review is warranted.
5 Scope and Limitations
EvoXplain does not assert that the clusters it identifies correspond to true causal mechanisms, nor that particular centroids represent any form of ground truth. Its outputs are contingent upon the attribution method used, and the observed structure may also reflect noise or inductive biases inherent to the model or the explanation technique. EvoXplain does not attempt to resolve these questions. Rather, its purpose is to bring such issues to the forefront and to encourage more cautious interpretation of explanation outputs.
EvoXplain has been designed to assess how models behave across repeated training runs. It determines whether a model settles into a single cluster of explanations or whether different explanations emerge across retraining. This perspective comes with several limitations.
First, while this work applies EvoXplain using attribution-based explanations, the framework itself is more general. Attribution vectors represent one way of operationalising interpretability, but other classes of explanations—such as causal abstractions, counterfactuals, symbolic rules, or circuit-level features—may also be suitable, provided they yield representations that can be quantified across multiple runs. As such, our findings instantiate using a particular family of explanation paradigms; extending EvoXplain to other forms of explanation is a natural direction for future work.
Second, EvoXplain detects the presence of multiple explanatory basins but does not assign any of them the status of a “correct” basin. Mechanism selection, validation, or prioritisation requires domain knowledge, causal evidence, and, in many settings, regulatory judgement. EvoXplain is intended to support such processes, not replace them.
Third, our empirical analysis is limited to tabular datasets and a small number of widely used model classes. While these settings are well studied and practically relevant, EvoXplain has not yet been evaluated across all data modalities, model architectures, or explanation techniques.
Fourth, EvoXplain assumes that repeated training is feasible. For many tabular and classical machine learning models this assumption is easily satisfied. For large-scale models, such as foundation systems, full retraining may be impractical, and alternative strategies—such as checkpoint sampling or multiple fine-tuning runs—may be required.
These limitations do not diminish the framework, but rather clarify the scope of its current applicability. EvoXplain is a diagnostic tool: it does not provide definitive answers, but instead surfaces stability issues that are often overlooked. Future work may extend the framework to additional explanation paradigms, model classes, data types, causal validation strategies, and decision criteria for interpreting and resolving basin structure.
6 Experimental Setup
6.1 Datasets
We evaluate EvoXplain on two tabular binary classification datasets.
Breast Cancer Wisconsin (Diagnostic): , . The task is to classify tumors as malignant or benign. Data are loaded from sklearn.datasets.load_breast_cancer.
COMPAS Recidivism: . The task is to predict two-year recidivism. We use the ProPublica COMPAS dataset with standard filtering (screening window, valid charge degree, non-missing scores). Categorical variables are one-hot encoded.
For each dataset, we perform stratified train–test splits with test fraction and fixed split seeds. All analyses are repeated independently per split.
6.2 Models
We evaluate the following model classes:
Logistic Regression (LogReg): -regularised logistic regression with solvers lbfgs (deterministic).
Random Forest (RF): Bootstrap aggregation with varying tree hyperparameters.
Linear models are trained on standardised features using StandardScaler. Tree-based models are trained on raw features.
6.3 Repeated Training Protocol
For each of the 10 dataset–model–split combinations (with five splits used for Random Forests), we perform independent training runs.
Each run differs only in:
the random seed used for optimisation and internal sampling,
the regularisation or model hyperparameters when explicitly varied.
For Logistic Regression, the regularisation strength is either:
fixed (), or
varied deterministically over a log-spaced grid between and .
For Random Forests, hyperparameters are sampled from predefined grids over: number of estimators, maximum depth, minimum samples per leaf, and maximum features.
Each training run produces a trained model, test accuracy, and a global explanation vector. Runs are executed in parallel on HPC infrastructure using chunked execution and aggregated post hoc (See the code repository). Due to the computational cost of Random Forest explanations, RF experiments were conducted over five independent train–test splits, which were sufficient to characterise the qualitative geometry of the explanation landscape.
Although the lbfgs solver is deterministic for fixed data and fixed hyperparameters, explanation multiplicity arises because training spans a range of regularisation strengths, allowing optimisation to settle into different equally optimal basins across the hyperparameter-conditioned loss landscape. Thus, the stochasticity measured by EvoXplain reflects basin selection under admissible training configurations, not randomness in optimisation or attribution.
6.4 Explanation Generation
For each trained model, we compute global feature importance using SHAP.
Logistic Regression: shap.LinearExplainer
Random Forest: shap.TreeExplainer
SHAP values are computed on the held-out test set. For each run , we obtain a single global importance vector
by averaging absolute SHAP values across test samples and outputs.
6.5 Aggregation and Clustering
For each split, explanation vectors are aggregated across runs and reordered using run indices to ensure completeness.
Explanation vectors are -normalised prior to clustering. Degenerate cases (nearly identical vectors) are detected using a fixed tolerance.
Clustering is performed using -means with candidate values
The optimal number of clusters is selected by maximising the silhouette score. Cluster centroids, sizes, and mechanistic entropy are computed as described in 4.3.
6.6 Disagreement Inspection
When , we identify representative runs for each explanatory basin by selecting the run closest to the cluster centroid in cosine distance.
Representative models are retrained and evaluated on the same test set. For each test instance , we compute the disagreement score
where denotes a representative model from basin .
Instances with the largest are reported to characterise cases where distinct explanatory modes lead to materially different predictions.
6.7 Visualisation
For each split, explanation vectors are projected into two dimensions using PCA after centring. Per-split manifolds display cluster structure and regularisation gradients.
A universal manifold is constructed by stacking explanation vectors across all splits and applying PCA without per-split centring, enabling comparison of global geometric structure across data realisations.
7 Results
7.1 Logistic Regression: explanation multiplicity under stochastic training
7.1.1 Breast Cancer
We begin with Logistic Regression on the Breast Cancer dataset, the simplest model in our study and, in principle, one where explanations should be stable. When regularisation strength is allowed to vary, this expectation does not hold.
For a representative split (split seed 109), we analyse 1,000 independent training runs. In explanation space, the SHAP vectors do not form a single cloud. Instead, they organise along a smooth, low-dimensional arc with clear separation between regions (Fig. 1). Clustering in explanation space yields an optimal solution at , with a silhouette score of 0.776, indicating well-separated explanation modes. The normalised explanation entropy is 0.922, showing that runs are distributed across the two modes rather than collapsing into a single dominant explanation.
The two clusters are not small perturbations of one another. Cluster 0 contains 663 runs with mean , while Cluster 1 contains 337 runs with mean . Despite occupying different regions of the explanation manifold, both clusters achieve comparable predictive accuracy (approximately 0.96). The accuracy distributions overlap substantially (Fig. 1, right panel). In other words, both explanation modes achieve similarly high predictive performance.
The clusters correspond to distinct regularisation regimes. The right panel of Fig. 1 shows a smooth gradient of values along the arc: strong regularisation (low ) concentrates at one end, while weak regularisation (high ) occupies the other. This gradient structure indicates that explanation geometry is tightly coupled to the regularisation–Loss landscape interaction.
These differences are reflected in the explanations themselves. The dominant features differ qualitatively between clusters. One mode emphasises classical size-related measurements, while the other places substantially more weight on variability- and shape-related features. The model is not merely reweighting the same features slightly differently; it is relying on qualitatively different explanatory patterns.
Importantly, the average attribution vector—computed as the arithmetic mean of all SHAP vectors across runs—does not lie on the populated manifold. In the split-level embedding, it falls in a sparsely populated region between clusters (Fig. 1), indicating that this average explanation does not correspond to a typical trained model instance.
When explanation vectors are stacked across all ten splits, the same qualitative picture persists. The universal explanation manifold shows ten parallel arcs, each corresponding to a different train–test realisation, that remain separated in explanation space, while predictive accuracy remains uniformly high along all tracks (Fig. 2). The global average attribution vector, computed across all runs and all splits, lies again away from the high-density regions of the manifold. This reinforces that averaging explanations across stochastic runs produces a vector that is not representative of any deployed model instance.
To assess whether these explanation differences translate into different predictions at the individual level, we compare representative models from the two clusters in split 109. In all held test instances, the mean absolute disagreement in the predicted malignant probability is 0.067 and the median is 0.035. However, the tail of the distribution is substantial: the maximum disagreement reaches , with 34 instances exceeding , 12 instances exceeding , and 4 instances exceeding . For the highest-disagreement case, one representative model assigns a malignant probability substantially different from the other, despite both models achieving high overall accuracy. These discrepancies arise even though aggregate performance metrics are indistinguishable.
Control Experiment: Stability under Fixed Regularisation. As a control, we repeat the same analysis under fixed regularisation strength (). Across ten independent train–test splits, each with 1,000 runs, the explanation vectors consistently collapse to a single cluster (), with negligible dispersion in explanation space. This confirms that the observed explanation multiplicity is not an artefact of the SHAP method or dimensionality reduction, but emerges specifically when the training process admits multiple equally accurate models.
| Split | Runs | Silh. () | Acc. mean | Acc. std | ||
| 100 | 1000 | 2 | 0.771 | 0.919 | 0.974 | 0.004 |
| 101 | 1000 | 2 | 0.795 | 0.943 | 0.961 | 0.007 |
| 102 | 1000 | 2 | 0.757 | 0.925 | 0.971 | 0.010 |
| 103 | 1000 | 2 | 0.774 | 0.932 | 0.976 | 0.012 |
| 104 | 1000 | 2 | 0.770 | 0.929 | 0.968 | 0.010 |
| 105 | 1000 | 2 | 0.806 | 0.944 | 0.956 | 0.005 |
| 106 | 1000 | 2 | 0.817 | 0.912 | 0.967 | 0.009 |
| 107 | 1000 | 2 | 0.761 | 0.920 | 0.957 | 0.005 |
| 108 | 1000 | 2 | 0.791 | 0.938 | 0.969 | 0.008 |
| 109 | 1000 | 2 | 0.776 | 0.922 | 0.962 | 0.007 |
| Split | Cl. | Runs | mean | std | Acc. mean | Acc. std | |
| 100 | 0 | 666 | 10.139 | 19.776 | [0.010, 99.4] | 0.974 | 0.004 |
| 100 | 1 | 334 | 10.110 | 19.318 | [0.010, 96.3] | 0.974 | 0.004 |
| 101 | 0 | 640 | 10.011 | 19.784 | [0.010, 99.9] | 0.961 | 0.007 |
| 101 | 1 | 360 | 14.598 | 23.484 | [0.010, 95.0] | 0.960 | 0.007 |
| 102 | 0 | 660 | 11.190 | 20.888 | [0.010, 99.6] | 0.971 | 0.010 |
| 102 | 1 | 340 | 11.889 | 20.636 | [0.010, 99.1] | 0.971 | 0.009 |
| 103 | 0 | 652 | 10.487 | 20.645 | [0.010, 99.3] | 0.976 | 0.012 |
| 103 | 1 | 348 | 12.280 | 21.910 | [0.011, 99.0] | 0.976 | 0.011 |
| 104 | 0 | 656 | 11.389 | 21.649 | [0.010, 99.2] | 0.969 | 0.010 |
| 104 | 1 | 344 | 11.722 | 20.984 | [0.010, 99.8] | 0.968 | 0.010 |
| 105 | 0 | 638 | 10.422 | 20.393 | [0.010, 97.1] | 0.956 | 0.005 |
| 105 | 1 | 362 | 11.910 | 20.600 | [0.010, 92.1] | 0.955 | 0.005 |
| 106 | 0 | 673 | 10.826 | 20.983 | [0.010, 98.9] | 0.967 | 0.009 |
| 106 | 1 | 327 | 10.000 | 19.920 | [0.010, 97.8] | 0.968 | 0.009 |
| 107 | 0 | 665 | 10.796 | 19.679 | [0.010, 97.2] | 0.958 | 0.005 |
| 107 | 1 | 335 | 11.651 | 21.597 | [0.011, 94.0] | 0.957 | 0.005 |
| 108 | 0 | 645 | 10.157 | 20.061 | [0.010, 98.6] | 0.970 | 0.008 |
| 108 | 1 | 355 | 10.767 | 20.794 | [0.010, 99.7] | 0.969 | 0.008 |
| 109 | 0 | 663 | 10.844 | 20.565 | [0.010, 99.7] | 0.962 | 0.007 |
| 109 | 1 | 337 | 12.199 | 21.267 | [0.011, 98.3] | 0.962 | 0.007 |
Across splits, Logistic Regression on the Breast Cancer dataset consistently exhibits two well-separated explanation basins under stochastic training (Table 1). In every split, clustering selects explanation modes with high silhouette scores (0.76–0.82) and near-maximal normalised entropy (0.91–0.94), indicating that runs are distributed across multiple explanatory mechanisms rather than collapsing into a single dominant solution. This structure is stable across independent train–test realisations.
Importantly, predictive performance remains uniformly high across splits and clusters. Mean accuracy ranges from 0.956 to 0.976, with small standard deviations, and accuracy distributions overlap substantially between explanation basins within each split. There is no split in which explanation multiplicity coincides with degraded or unstable predictive performance, consistent with the split-level and universal explanation manifolds.
Cluster-level analysis (Table 2) further shows that the two explanation basins are not reducible to a simple partition by regularisation strength. Within each split, both clusters span the full sampled range of values, with similar mean and variance. This indicates that the mechanistic separation observed in explanation space reflects genuinely distinct explanatory modes rather than trivial hyperparameter binning. Note that while position along the arc correlates smoothly with regularisation strength (Fig.1, right panel), cluster membership does not partition by C: runs with identical regularisation can land in different basins depending on other aspects of the optimisation trajectory. As a control, when regularisation strength is fixed, explanations collapse to a single cluster across all splits, confirming that explanation multiplicity emerges specifically when the training process admits multiple equally accurate solutions.
7.1.2 COMPAS
Results for Logistic Regression on the COMPAS dataset are summarised in Figures 3 and 4 and in Tables 3 and 4. Across all ten train–test splits, explanation clustering consistently selects two clusters (), with silhouette scores ranging from 0.73 to 0.91 (Table 3). These values indicate strong and stable separation between explanation modes across resampled datasets—notably higher than those observed on Breast Cancer. The normalised entropy ranges from 0.83 to 0.91, indicating that explanation modes are somewhat unequally populated, with cluster sizes typically in a ratio of approximately 1:2.5.
Predictive performance remains stable across splits. Mean accuracy ranges from 0.68 to 0.70 (this is typical for recidivism prediction tasks given inherent unpredictability), with standard deviations on the order of or smaller (Table 3). There is no evidence that the emergence of multiple explanation modes is associated with unstable or degraded performance. This is consistent with the split-level manifolds (Fig. 3), where accuracy varies smoothly along the explanation geometry rather than separating sharply by cluster.
Table 4 shows that, as with Breast Cancer, both explanation clusters span the full range of regularisation strengths. In every split, both clusters contain runs with values ranging from 0.01 to approximately 100, with similar mean values of around 10–12. This indicates that the mechanistic separation in explanation space is not a simple partition by regularisation strength. Rather, the two explanation modes represent genuinely distinct regions of the explanation landscape that models can reach regardless of their regularisation setting. Predictive accuracy remains effectively identical across clusters within each split, with differences typically below 0.001.
When explanation vectors are stacked across all splits, the same structure persists (Fig. 4). The universal explanation manifold consists of multiple separated tracks that remain distinct across resampled datasets, while accuracy remains confined to a narrow band. As in the split-level analysis, the global average attribution vector lies away from the populated regions of the manifold, indicating that averaging explanations across stochastic runs does not yield a representative explanation.
Control Experiment: Stability under Fixed Regularisation. As a control, we repeat the same analysis under fixed regularisation strength (). Across all ten independent train–test splits, each with 1,000 runs, the explanation vectors consistently collapse to a single cluster (), with zero entropy and negligible dispersion in explanation space. This confirms that the observed explanation multiplicity is not an artefact of the SHAP method or clustering procedure, but emerges specifically when the training process admits multiple equally accurate models through varied regularisation.
Therefore, the COMPAS results reinforce the findings from Breast Cancer: explanation multiplicity in Logistic Regression is highly consistent across splits, while predictive performance remains effectively unchanged. Importantly, the mechanistic separation is not driven by hyperparameter differences—models with similar values can land in different explanation basins, indicating that multiplicity reflects the structure of the explanation landscape itself.
| Split | Runs | Silh. () | Acc. mean | Acc. std | ||
| 100 | 1000 | 2 | 0.802 | 0.834 | 0.693 | 0.001 |
| 101 | 1000 | 2 | 0.910 | 0.841 | 0.695 | 0.002 |
| 102 | 1000 | 2 | 0.743 | 0.908 | 0.691 | 0.001 |
| 103 | 1000 | 2 | 0.814 | 0.846 | 0.684 | 0.001 |
| 104 | 1000 | 2 | 0.803 | 0.870 | 0.689 | 0.003 |
| 105 | 1000 | 2 | 0.833 | 0.897 | 0.688 | 0.001 |
| 106 | 1000 | 2 | 0.767 | 0.897 | 0.682 | 0.001 |
| 107 | 1000 | 2 | 0.873 | 0.855 | 0.685 | 0.001 |
| 108 | 1000 | 2 | 0.841 | 0.855 | 0.684 | 0.002 |
| 109 | 1000 | 2 | 0.727 | 0.844 | 0.698 | 0.001 |
| Split | Cl. | Runs | mean | std | Acc. mean | Acc. std | |
| 100 | 0 | 735 | 9.338 | 18.396 | [0.010, 99.4] | 0.693 | 0.001 |
| 100 | 1 | 265 | 12.323 | 22.541 | [0.010, 99.2] | 0.693 | 0.001 |
| 101 | 0 | 730 | 12.290 | 21.532 | [0.010, 98.8] | 0.695 | 0.001 |
| 101 | 1 | 270 | 9.964 | 20.581 | [0.010, 99.9] | 0.695 | 0.002 |
| 102 | 0 | 677 | 11.069 | 20.082 | [0.010, 99.6] | 0.691 | 0.001 |
| 102 | 1 | 323 | 12.180 | 22.227 | [0.010, 99.5] | 0.691 | 0.001 |
| 103 | 0 | 727 | 11.190 | 20.780 | [0.010, 99.0] | 0.684 | 0.001 |
| 103 | 1 | 273 | 10.901 | 21.966 | [0.010, 99.3] | 0.684 | 0.001 |
| 104 | 0 | 709 | 11.658 | 21.687 | [0.010, 99.8] | 0.689 | 0.003 |
| 104 | 1 | 291 | 11.129 | 20.762 | [0.010, 92.7] | 0.689 | 0.003 |
| 105 | 0 | 687 | 11.334 | 20.312 | [0.010, 97.1] | 0.688 | 0.001 |
| 105 | 1 | 313 | 10.141 | 20.824 | [0.010, 96.9] | 0.688 | 0.001 |
| 106 | 0 | 687 | 10.592 | 21.029 | [0.010, 98.9] | 0.682 | 0.001 |
| 106 | 1 | 313 | 10.476 | 19.777 | [0.010, 97.4] | 0.682 | 0.001 |
| 107 | 0 | 720 | 11.347 | 20.892 | [0.011, 97.2] | 0.685 | 0.001 |
| 107 | 1 | 280 | 10.400 | 18.852 | [0.010, 89.9] | 0.685 | 0.001 |
| 108 | 0 | 720 | 11.047 | 21.038 | [0.010, 99.7] | 0.684 | 0.002 |
| 108 | 1 | 280 | 8.642 | 18.258 | [0.011, 93.3] | 0.685 | 0.002 |
| 109 | 0 | 728 | 11.576 | 20.999 | [0.010, 98.3] | 0.698 | 0.001 |
| 109 | 1 | 272 | 10.562 | 20.295 | [0.012, 99.7] | 0.698 | 0.001 |
7.2 Disagreement Between Equally Accurate Explanation Basins
To assess whether explanation multiplicity can translate into materially different predictions at the individual level, we compare representative models drawn from distinct explanation basins within the same train–test split. For each held-out test instance , we compute the absolute difference in predicted malignant probability,
where and denote representative models selected from two different explanation clusters. Importantly, the compared models share the same model class, training data, and optimisation objective, and achieve comparable test accuracy. This controls for performance differences and isolates disagreement attributable to distinct explanatory mechanisms rather than optimisation noise or underfitting.
On the Breast Cancer dataset, disagreement magnitudes span a wide range. While the mean disagreement across test instances is typically modest, the upper tail is substantial. Figure 5 illustrates this phenomenon for a representative split (Split 107), in which 76 test instances exhibit disagreement exceeding , 33 exceed , and 15 exceed . The maximum observed disagreement in this split reaches , despite both representative models achieving similar overall accuracy (approximately 0.95). These values indicate that explanation multiplicity can manifest as large probability divergences for individual patients even when aggregate performance metrics are indistinguishable.
The highest-disagreement case in Figure 5 corresponds to a malignant tumour for which one model assigns a probability of 0.72 while the other assigns 0.04, leading to opposite clinical decisions for the same patient. Notably, this disagreement does not arise from marginal predictions near the decision threshold: one model is confidently malignant, while the other is confidently benign. The accompanying SHAP attributions reveal qualitatively different patterns of feature reliance across the two explanation basins. Although both models attend to overlapping morphological feature families, they weight and combine this evidence differently, resulting in incompatible diagnostic conclusions. This behaviour is inconsistent with an interpretation of disagreement as stochastic fluctuation or local instability, and instead reflects the presence of distinct explanatory mechanisms that remain viable at comparable accuracy.
Across all ten train–test splits, the magnitude of individual-level disagreement varies substantially. Maximum disagreement per split ranges from negligible values (e.g., in split 103) to large discrepancies exceeding (e.g., splits 101 and 107). This variability indicates that the practical manifestation of explanation multiplicity depends on the particular data realisation. However, critically, large individual-level disagreements can arise even when aggregate accuracy and calibration metrics remain effectively unchanged. In such cases, standard performance summaries obscure the fact that different model instances may assign substantially different risk estimates to the same individual.
Crucially, disagreement is not confined to cases where models predict different labels. Table 5 reports representative test instances from Split 107 for which both explanation basins correctly predict the same class, yet assign substantially different malignant probabilities (). In these cases, predictions agree at the label level, but differ meaningfully in confidence, with one model indicating borderline risk and the other indicating high certainty. While these instances do not lead to immediate label flips, they demonstrate that mechanistic disagreement can persist even when classification outcomes are identical, manifesting as divergent belief states rather than discrete decision changes. Such differences are relevant in downstream clinical contexts where predicted probabilities inform risk stratification, follow-up decisions, or human oversight.
Taken together, these results show that explanation multiplicity is not merely a geometric or cosmetic property of explanation space. When multiple explanation basins coexist at similar accuracy, the choice of model instance can materially affect individual-level predictions, either by altering the predicted class or by substantially shifting estimated risk. The observed clusters therefore correspond to qualitatively distinct explanatory mechanisms, not small perturbations of a single underlying solution. High predictive performance alone does not imply mechanistic stability or explanatory uniqueness.
To further characterise the structure of these mechanisms, we examine the characteristic explanation patterns associated with each basin. Figure 5 visualises the centroid attribution profiles of the explanatory modes identified on the Breast Cancer dataset, highlighting how equally accurate models systematically rely on different feature groupings when forming predictions.
| Test index | True label | |||
| 36 | Benign | 0.79 | 1.00 | 0.21 |
| 169 | Malignant | 0.60 | 0.81 | 0.21 |
7.3 Random Forests: explanation multiplicity in ensemble models
Having established that explanation multiplicity arises in Logistic Regression under stochastic training—despite its convex objective and stable predictive performance—we next examine whether the same phenomenon appears in a qualitatively different model class. Random Forests introduce non-convexity, ensemble averaging, and additional sources of stochasticity through bootstrap sampling and feature subsampling. Evaluating explanation stability in this setting allows us to test whether the observed multiplicity is specific to linear models under hyperparameter variation, or whether it persists in more expressive and widely deployed ensemble architectures.
7.3.1 Breast Cancer
Results for Random Forests on the Breast Cancer dataset are shown in Fig. 6 and Fig. 7. Each train–test split aggregates approximately Random Forest instances, reflecting stochasticity arising from bootstrap resampling, feature subsampling, and hyperparameter variation. This large-scale sampling allows the structure of explanation space to be examined at high resolution.
Figure 6 visualises the universal explanation manifold obtained by stacking explanation vectors across multiple train–test splits. In contrast to the smooth, curved manifolds observed for Logistic Regression, Random Forest explanation vectors organise into multiple dense and well-separated regions. These regions persist across splits, indicating stable explanation multiplicity in the ensemble setting. As in the linear case, the global average attribution vector—computed by averaging SHAP values across all runs—lies outside the high-density regions of the manifold. This indicates that aggregating explanations across stochastic training runs again produces a representation that does not correspond to a typical deployed model instance.
Importantly, predictive accuracy varies smoothly across the explanation manifold and does not align with any single cluster or region. Models drawn from different explanation clusters achieve comparable test accuracy, with no sharp performance boundary separating the modes. This demonstrates that, in Random Forests, explanation multiplicity is not driven by accuracy trade-offs or poorly performing solutions, but instead reflects the coexistence of distinct explanatory mechanisms within the same performance regime.
Figure 7 shows the explanation landscape for a single representative split (Split 104). Even within a fixed train–test realisation, explanation vectors again separate into two high-density regions. The first principal component captures the majority of variance in explanation space, while the second component reflects residual spread within each mode. As in the universal manifold, predictive accuracy varies continuously across both regions, indicating that explanation multiplicity arises intrinsically within a single split and does not depend on aggregation across data resamplings or extreme hyperparameter configurations.
Taken together, these results show that explanation multiplicity persists in Random Forests despite ensemble averaging and strong inductive bias towards prediction stability. In contrast to Logistic Regression, Random Forests often exhibit limited disagreement in predicted labels or probabilities at the individual level, reflecting the stabilising effect of ensembling. However, this output stability masks substantial mechanistic non-uniqueness: multiple, distinct explanation basins coexist at comparable accuracy. Thus, even when predictions appear stable, the underlying reasoning employed by the model remains non-unique.
This contrast highlights an important distinction between predictive stability and mechanistic stability. Ensemble methods may suppress variability in outputs while preserving—or even amplifying—multiplicity in explanation space. High predictive accuracy and low output variance therefore do not imply a unique or well-defined explanatory mechanism.
Preliminary experiments on additional datasets and model classes (including Adult Income and Elastic Nets) suggest similar patterns of explanation multiplicity. These analyses are not reported here and are provided in the accompanying code repository as exploratory results.
8 Discussion
EvoXplain shows that even models with high prediction accuracy do not guarantee explanation stability, even when datasets are thought to be ”simple”. Across both datasets, multiple model classes resulted in multiple explanatory modes with considerable support. This suggests that interpretability cannot be inferred from model form or accuracy alone. For example, linear models commonly assumed to be transparent do not consistently converge to a single explanatory mechanism under standard retraining procedures. In our experiments, Logistic Regression exhibits explanation multiplicity across datasets when regularisation is allowed to vary, despite high and stable predictive accuracy (Lipton (2018); Rudin (2019)). This challenges the common practice of treating transparency and accuracy as sufficient conditions for interpretability. The goal is not to eliminate variability, but to characterise the range of mechanisms that arise under standard retraining procedures commonly used in practice.
A dominant operational assumption in applied machine learning is that inspecting a single trained model instance is sufficient to characterise a system’s reasoning. EvoXplain clarifies the limits of that assumption, as it shows that models with the same accuracy can have different internal reasoning structures. This implies that one attribution profile is likely just one out of many possible ones. Accordingly, interpretability should be framed as a population-level property of trained models rather than as an attribute of a single fitted instance. While predictive performance remains stable, explanation multiplicity matters because explanations are often used to justify, audit, or regulate model-driven decisions at the individual level. This suggests that, rather than just focusing on a single model instance, we need to develop stability-aware interpretability of models that provides explanations based on the behaviour of models over multiple instances or runs (Arrieta and others (2020); Selbst and Barocas (2018)).
EvoXplain does not claim to identify where multiplicity originates—whether from data structure, model flexibility, or optimisation randomness. Instead, it makes this uncertainty visible. The framework is deliberately agnostic to causal origin and instead focuses on empirical detectability.
The multiplicity observed cannot be attributed to instability or noise in SHAP itself. SHAP values are deterministic for any fixed model. EvoXplain examines attribution profiles across independent model trainings, rather than variability within a single explanation. This distinction is critical: EvoXplain characterises variability across trained models, not variability within an attribution method. Because SHAP is deterministic for a fixed model, any observed clustering reflects differences between trained models rather than stochasticity in the explainer itself. As such, the resulting clusters reflect real differences in learned mechanisms. For the Breast Cancer benchmark, the centroids of the discovered basins reinforce established diagnostic attributes such as perimeter, radius, area, concavity, and associated morphological descriptors used in clinical tumour assessment. The emergence of distinct, clinically meaningful feature profiles across basins would not be expected if clustering were driven by random noise.
Therefore, inconsistency between explanations is not something to be averaged out as noise. Instead, we reframe it as an empirical observation. In this sense, explanation multiplicity is a property to be measured, reported, and reasoned about, rather than suppressed. In the presence of multiple modes of explanation, the decision to select one becomes a question of science or governance, rather than something the training pipeline can resolve automatically (European Commission (2023); Rudin (2019)).
This perspective aligns with recent safety assurance methodologies for machine learning, such as AMLAS and the “BIG Argument” for AI safety proposed by Habli and colleagues (Hawkins et al. (2021); Habli et al. (2025)), which emphasise the need for explicit evidence of model robustness beyond predictive accuracy alone. Within such frameworks, explanation stability is not merely an interpretability preference but a form of verification evidence. EvoXplain contributes precisely this type of evidence by quantifying how model reasoning varies under standard retraining procedures, thereby supporting or challenging claims of reliability in safety-critical ML systems (Paterson et al. (2025)).
A closely related operational response to model multiplicity is to aggregate explanations over a set of near-optimal models and report uncertainty around the aggregate. For instance, Rashomon-set approaches in AutoML construct a set of models within a performance tolerance and then average partial dependence profiles across that set to obtain a single ”Rashomon PDP”, optionally accompanied by confidence intervals to visualise disagreement Cavus et al. (2025). EvoXplain highlights a stronger failure mode: when explanation variability forms separable basins of near-equivalent accuracy, aggregation is not merely lossy but can be methodologically misleading, producing a ghost explanation that does not correspond to any model in the population. In such settings, interpretability requires representing the set of mechanistic regimes and their relative support (frequency), rather than collapsing them into a single averaged explanation with uncertainty bands.
This is particularly important because explanations have immediate consequences in certain domains. In criminal justice contexts, multiple basins of attraction counteract the assumption that a model provides a singular, stable justification Umbach et al. (2023). In biomedicine, instability observed even on standard datasets warns that attribution-based interpretations should be treated with caution when used to support clinical claims Agrawal and Jain (2023). Thus, EvoXplain helps to identify cases where a single explanation is insufficient to account for the behaviour of a model under retraining. In both settings, explanation stability becomes relevant precisely because explanations are used as part of decision-making processes rather than as post-hoc descriptions.
In our experiments, we focus on tabular learning, where repeated training is manageable. However, retraining large language models (LLMs) presents a more difficult challenge. We do not claim that the specific geometric structures observed here generalise directly to large-scale models; rather, we argue that the underlying issue of mechanistic non-identifiability is likely to be amplified in such settings. LLMs train in very high-dimensional, non-convex spaces, are exposed to stochastic data, and follow fine-tuning paths that may diverge substantially. Whether models that exhibit similar external behaviour in such settings rely on the same internal mechanisms remains an open empirical question. Measuring this at scale is challenging, but not impossible: models and their fine-tuned variants can be compared, checkpoints sampled, and attribution patterns analysed across independent runs (Olah et al. (2020)). EvoXplain provides a conceptual starting point. The instability observed even in small models suggests that non-identifiability in foundation models may be important and deserves further study (Lakshminarayanan et al. (2017)).
Beyond the empirical evidence, mechanistic multiplicity underscores the need to evaluate reliability and interpretability as properties that cannot be assumed. Instead, they must be empirically established under realistic training variability.
Why Mechanistic Multiplicity Matters
Mechanistic multiplicity primarily raises questions of reliability, rather than philosophy. It concerns whether a modelling pipeline produces consistent reasoning under retraining, not whether a single explanation appears plausible. When predictive accuracy decays, retraining is standard practice, and the modelling process may generate different reasoning outputs, with no guarantee of stable reasoning over time. Reliability must therefore concern the process rather than a particular model instance. In high-stakes contexts, explanations justify actions. Even at equal predictive accuracy, models can diverge in the actions, risk factors, and scientific hypotheses they imply, settling into different basins of explanation. This divergence is not resolved by appeals to sparsity, simplicity, or accuracy alone. When explanations inform decisions rather than merely describe them, instability in the underlying mechanism becomes a substantive safety and interpretability issue. Appeals to simplicity do not resolve multiplicity. In the presence of competing explanation basins with equivalent sparsity, complexity, and predictive accuracy, Occam’s razor cannot favour one over another. What remains is a form of empirical underdetermination. In many cases, competing explanation basins are comparable not only in accuracy, but also in sparsity and apparent simplicity, leaving no principled criterion for selection based solely on model-internal properties. In such cases, the presence of multiple stable explanatory modes evidences a profound underdetermination, which EvoXplain is designed to identify and document.
There are boundaries to this work. We analyse explanation stability using attribution methods; other interpretability approaches, such as causal abstraction or circuit tracing, may offer different perspectives (Olah et al. (2020)). EvoXplain does not attempt to correct bias, determine which basin is “correct”, or quantify downstream risks such as fairness or failure modes Selbst and Barocas (2018). We consider only a limited number of datasets and model classes; this study represents an initial step toward broader analysis across architectures, attribution schemes, and domains such as reinforcement learning or language modelling. These limitations define the scope of the present study rather than its validity. The contribution of EvoXplain is not to propose a new explainer, but to provide a diagnostic framework that makes explanation instability measurable, comparable, and reportable across model populations.
Despite these limitations, EvoXplain opens new research questions. Are explanation basins correlated with latent subgroups, causal mechanisms, or inductive bias pathways (Marx et al. (2020))? Is it possible to prefer certain basins for safety, fairness, or generalisability (Rudin (2019))? More generally, EvoXplain shifts interpretability from analysing individual models to analysing populations of models trained under the same procedure. This reframing allows explanation stability to be measured, compared, and monitored at the level of model populations.
To summarise, EvoXplain treats interpretability as a property of model behaviour under retraining. By focusing on mechanistic multiplicity rather than assuming a single explanatory mode, EvoXplain provides tools to uncover and characterise multiplicity when it arises. This, in turn, raises methodological, epistemic, and governance challenges. If models admit multiple internally consistent mechanistic explanations, then explanation selection, validation, and governance require explicit criteria. If models embody a multiplicity of mechanistic stories, how should they be assessed, described, or governed? While EvoXplain does not resolve these questions, it makes them explicit and positions explanation stability as a concrete target for future research. In practice, EvoXplain enables practitioners and auditors to detect when single-model explanations are insufficient, without prescribing which explanation should be preferred.
9 Conclusion
Explanations are often treated as fixed model properties—examined once and assumed to reflect one underlying mechanism. EvoXplain argues the opposite, and instead quantifies this assumed stability. Across two benchmarks and two model families, models achieved high predictive performance. However, their explanations frequently diverged into multiple distinct modes, even within the same model class and data split. The extent and structure of this multiplicity varied by dataset—for example, explanation basins were more evenly balanced on the Breast Cancer dataset and more asymmetric on COMPAS—but in all cases, high predictive accuracy did not imply a single stable mechanism. This demonstrates that explanation stability cannot be inferred from predictive performance, model class, or perceived simplicity alone.
EvoXplain offers a diagnostic view rather than a decision rule. It shows that multiplicity exists, shows its form, and specifies the frequency at which different explanatory mechanisms can emerge. It does not dictate which local minima should be prioritized; that is a decision to be made from the domain knowledge, validation or regulatory governance. It is a valuable tool to make multiplicity non-abstract and measurable, and to encourage interpretability practices that account for stability. In this sense, EvoXplain complements existing interpretability methods by exposing when single-model explanations are insufficient.
Future research will focus on exploring if mechanistic multiplicity is present within other paradigms and methods of training and attribution. The phenomenon is likely to be much broader than what is presented here, especially in areas where the decisions of the models have real consequences, like genomics and biomedical risk modelling. The question is further amplified within foundation models. Do models that function similarly, internally follow the same mechanism, or do they split into distinct explanatory modes at scale? Answering this question will require population-level analysis of trained models rather than inspection of isolated instances.
Interpretability can no longer centre on a single trained model. The assumed single explanation is viewed as a distribution of explanations, and their stability, diversity, and governance relevance are not assumptions, but empirical matters that require measurement. EvoXplain pushes the field in this direction by treating explanation stability as a primary diagnostic of model behaviour. By making explanation multiplicity observable and quantifiable, EvoXplain provides a foundation for future interpretability, auditing, and governance frameworks that operate at the level of model populations rather than individual fits.
Acknowledgements
The EvoXplain framework is the subject of a provisional patent filing. The experiments reported here were run on the University of Hertfordshire High Performance Computing Cluster, whose support is gratefully acknowledged.
Competing Interests
EvoXplain is the subject of a filed provisional patent application. The author declares no other competing interests.

### v3

@@
 We introduce EvoXplain, a simple diagnostic framework that measures the stability of model explanations across repeated training.
-Rather than analysing the explanation of a single trained model, EvoXplain treats explanations as samples drawn from the stochastic optimisation process itself—without aggregating predictions or constructing ensembles—and examines whether these samples form a single coherent explanation or instead separate into multiple, distinct explanatory modes.
+Rather than analysing the explanation of a single trained model, EvoXplain treats explanations as samples drawn from the training and model selection pipeline itself—without aggregating predictions or constructing ensembles—and examines whether these samples form a single coherent explanation or instead separate into structured regions of explanation space corresponding to distinct explanatory modes.
 We evaluate EvoXplain on the Breast Cancer and COMPAS datasets across two widely used model classes: Logistic Regression and tree-based models (Random Forests).
 Although all models achieve high predictive accuracy, their explanations frequently exhibit clear multimodality.
-Even models commonly assumed to be stable—such as Logistic Regression—can produce multiple, well-separated explanatory basins under repeated training on the same data split.
-Importantly, these differences are not explained by hyperparameter variation or simple performance trade-offs.
+Even models commonly assumed to be stable—such as Logistic Regression—can give rise to structured, low-dimensional explanation manifolds that separate into recurrent regions under repeated training on the same data split.
+Crucially, these distinct explanation modes coexist even at near-identical hyperparameter configurations, indicating genuine explanation non-identifiability rather than smooth sensitivity to regularisation strength.
 EvoXplain does not attempt to select a “correct” explanation.
@@
 Despite the fact that practitioners routinely retrain models during development and performance tuning, interpretability is typically assessed using a single trained instance—or by averaging explanations across runs—rather than by explicitly examining whether explanations themselves remain consistent under retraining.
-This practice implicitly assumes that high predictive accuracy corresponds to a single, stable way of reasoning (Lipton (2018); Rudin (2019)).
+This practice implicitly assumes that high predictive accuracy corresponds to a single, stable way of reasoning (Lipton, 2018; Rudin, 2019).
 There have been sustained challenges to this assumption in the literature.
-Perhaps the most well-known is the Rashomon effect, which observes that distinct models may achieve similar predictive performance while relying on different internal logics (Breiman (2001); Marx et al.
-(2020)).
-Related work on underspecification highlights that modern learning pipelines often admit many solutions that are indistinguishable in terms of accuracy, leaving optimisation procedures to select arbitrarily among them (D’Amour et al.
-(2022)).
-Other studies have shown that explanations can vary substantially when models are retrained, or when features are correlated or perturbed, including documented failures of attribution methods under retraining, feature overlap, or input transformations (Kindermans et al.
-(2017); Nogueira et al.
-(2018); Kumar et al.
-(2020); Frye et al.
-(2020)).
+Perhaps the most well-known is the Rashomon effect, which observes that distinct models may achieve similar predictive performance while relying on different internal logics (Breiman, 2001; Marx et al., 2020).
+Related work on underspecification highlights that modern learning pipelines often admit many solutions that are indistinguishable in terms of accuracy, leaving training and model selection procedures to select arbitrarily among them (D’Amour et al., 2022).
+Other studies have shown that explanations can vary substantially when models are retrained, or when features are correlated or perturbed, including documented failures of attribution methods under retraining, feature overlap, or input transformations (Kindermans et al., 2017; Nogueira et al., 2018; Kumar et al., 2020; Frye et al., 2020).
 Collectively, these results suggest that explanation stability cannot be assumed a priori.
@@
 To address this gap, we propose EvoXplain, a simple and easy-to-apply framework for assessing the stability of model explanations across repeated training.
-Rather than evaluating a single attribution vector, EvoXplain treats explanations as outcomes of stochastic optimisation and analyses them as a distribution induced by the training process itself.
+Rather than evaluating a single attribution vector, EvoXplain treats explanations as outcomes of repeated training and model selection and analyses them as a distribution induced by the training pipeline itself.
 By clustering explanation samples obtained from repeated instantiations of the same model class—without aggregating predictions or constructing ensembles—EvoXplain determines whether training consistently yields a single explanatory mode or instead produces multiple, distinct explanatory regimes.
 EvoXplain does not aim to decide which explanation is correct.
-Instead, it provides evidence about whether a model’s explanations are reliable or whether they depend sensitively on the particular optimisation trajectory taken during training.
-This perspective aligns with a growing view that interpretability should be evaluated empirically rather than assumed, a position increasingly reflected in discussions of responsible AI, model auditing, and governance (Arrieta and others (2020); Selbst and Barocas (2018); European Commission (2023)).
+Instead, it provides evidence about whether a model’s explanations are reliable or whether they depend sensitively on the particular training and model selection trajectory taken during deployment and retraining.
+This perspective aligns with a growing view that interpretability should be evaluated empirically rather than assumed, a position increasingly reflected in discussions of responsible AI, model auditing, and governance (Arrieta and others, 2020; Selbst and Barocas, 2018; European Commission, 2023).
 By making explanation stability measurable, EvoXplain enables practical questions such as: does this model tell a consistent story, or does its reasoning change across retraining?
@@
 Linear models and decision trees, for instance, are often tagged as “interpretable” because practitioners are able to read and comprehend the parameters or structures involved.
-However, some recent contributions to the field argue that readability cannot be equated with the models achieving high predictive accuracy (Lipton (2018)).
+However, some recent contributions to the field argue that readability cannot be equated with the models achieving high predictive accuracy (Lipton, 2018).
 EvoXplain follows the same direction by considering interpretability as something that can be measured and verified rather than assumed.
 2.2 Rashomon Effects and Predictive Multiplicity
-Our work relates to the Rashomon effect in that different models can arrive at the same answer and yet rely on totally different paths or reasoning (Breiman (2001); Semengue et al.
-(2019)).
-Some recent research considers predictive multiplicity under settings of fairness and of unconstrained risk (Umbach et al.
-(2023); Marx et al.
-(2020); Fisher et al.
-(2019)), showing that the same risk scores can be justified through different models (Umbach et al.
-(2023)).
+Our work relates to the Rashomon effect in that different models can arrive at the same answer and yet rely on totally different paths or reasoning (Breiman, 2001; Semengue et al., 2019).
+Some recent research considers predictive multiplicity under settings of fairness and of unconstrained risk (Umbach et al., 2023; Marx et al., 2020; Fisher et al., 2019), showing that the same risk scores can be justified through different models (Umbach et al., 2023).
 Unlike these approaches, EvoXplain studies multiplicity within the same model class, trained on the same data, to see whether retraining leads to different explanation patterns.
@@
 2.4 Feature Importance Instability
-Previous studies indicate that SHAP, LIME, and gradients as explanation methods can produce instability due to retraining, dataset perturbations, or correlated features (Kindermans et al.
-(2017); Nogueira et al.
-(2018); Kumar et al.
-(2020); Agarwal et al.
-(2021)).
+Previous studies indicate that SHAP, LIME, and gradients as explanation methods can produce instability due to retraining, dataset perturbations, or correlated features (Kindermans et al., 2017; Nogueira et al., 2018; Kumar et al., 2020; Agarwal et al., 2021).
 Most of these studies focus on instability in some specific values.
@@
 While such approaches usefully acknowledge that multiple near-optimal models exist, they typically treat multiplicity as uncertainty around a single aggregated explanation.
-EvoXplain studies a different phenomenon: structured, discrete explanatory modes (basins) among equally accurate models, and shows that explanation aggregation can yield a ghost explanation that may not correspond to any realised model instance.
+EvoXplain studies a different phenomenon: low-dimensional explanation manifolds with structured regions among equally accurate models, and shows that explanation aggregation can yield a ghost explanation that may not correspond to any realised model instance.
 In other words, rather than summarising variability around one explanation, EvoXplain detects and quantifies distinct mechanistic regimes within the same training pipeline.
 2.6 Mechanistic Interpretability in Deep Models
-The understanding of circuits and the internal functioning of large neural networks remains an ongoing area of research (Olah et al.
-(2020)).
+The understanding of circuits and the internal functioning of large neural networks remains an ongoing area of research (Olah et al., 2020).
 Specific techniques tend to rely on handcrafted analyses or other tools engineered for individual designs.
@@
 Soares et al.
-introduce Meta Subspace Analysis (MSA), which discovers exceptional subspaces and relates their metafeatures (e.g., distributional and dependency summaries) to model (mis)behavior via association rules (Soares et al.
-(2025)).
+introduce Meta Subspace Analysis (MSA), which discovers exceptional subspaces and relates their metafeatures (e.g., distributional and dependency summaries) to model (mis)behavior via association rules (Soares et al., 2025).
 This perspective is complementary to EvoXplain.
@@
 Beyond interpretability research, the safety assurance community has emphasised the need for explicit evidence about model robustness under retraining.
-Habli and colleagues propose structured assurance methodologies for machine learning components, such as AMLAS and the “BIG Argument” for AI safety (Habli et al.
-(2025); Hawkins et al.
-(2021)), which require that system behaviour be justified and verified beyond predictive accuracy alone.
+Habli and colleagues propose structured assurance methodologies for machine learning components, such as AMLAS and the “BIG Argument” for AI safety (Habli et al., 2025; Hawkins et al., 2021), which require that system behaviour be justified and verified beyond predictive accuracy alone.
 EvoXplain contributes to this assurance perspective by offering a concrete diagnostic for explanation stability.
@@
 This assumption is not always justified.
-Due to randomness in the training process, model inductive biases, and model flexibility, it is possible for multiple different explanations to achieve similar predictive performance.
+Due to randomness in the training process, model inductive biases, and the non-identifiability of admissible training configurations within a modelling pipeline, it is possible for multiple different explanations to achieve similar predictive performance.
 A first step toward trustworthy interpretability is therefore to verify whether explanatory multiplicity exists at all.
-Formally, consider a model class trained on a fixed dataset .
+Formally, consider a model class trained on a fixed dataset using a fixed training pipeline (including preprocessing, model family, and admissible hyperparameter configurations).
 Standard interpretability workflows examine a single trained instance and its corresponding attribution vector .
-In contrast, this work asks the following question: when is trained multiple times with independent random initializations, do the resulting attribution vectors converge to a single explanatory pattern, or do they separate into multiple distinct patterns?
-Let denote independent training runs on , and let denote the corresponding explanation vectors. EvoXplain analyses the empirical distribution of to determine:
+In contrast, this work asks the following question: when is instantiated multiple times under the same pipeline—via different random seeds, bootstrap samples, or admissible hyperparameter choices—do the resulting attribution vectors converge to a single explanatory pattern, or do they separate into multiple distinct patterns?
+Let denote independent model instances produced by this pipeline on , and let denote the corresponding explanation vectors. EvoXplain analyses the empirical distribution of to determine:
 1.
@@
 In practice, we repeat this analysis across multiple train–test splits of to assess the robustness of explanatory behaviour under data resampling.
-However, the core object of interest remains the distribution of explanations induced by repeated training within a fixed data realisation.
+However, the core object of interest remains the distribution of explanations induced by repeated admissible instantiations of the same training pipeline on a fixed data realisation.
 EvoXplain offers an empirical method for answering these questions.
-Rather than assuming explanations are stable, it treats them as random samples generated by the training process and assesses their structure using clustering and entropy.
-This reframes interpretability as an empirical question: whether a model class tends to produce a single coherent explanation or multiple competing explanatory mechanisms under repeated instantiation.
+Rather than assuming explanations are stable, it treats them as samples generated by the training pipeline and assesses their structure using clustering and entropy.
+This reframes interpretability as an empirical question: whether a model class, under realistic retraining and model selection, tends to produce a single coherent explanation or multiple competing explanatory mechanisms.
 Contributions
@@
 Instead of assuming that a model has a single reasoning way, we think of explanations as samples that can be analysed over several training runs.
-Second, we introduce EvoXplain, a simple diagnostic framework that determines if explanations cluster into a single pattern or if they are dispersed across multiple clusters of different patterns.
+Second, we introduce EvoXplain, a simple diagnostic framework that determines if explanations cluster into a single pattern or if they are dispersed across multiple basins of different patterns.
 It also measures the occurrence frequency of each distinct explanatory pattern.
@@
 Let denote a model class (e.g., Logistic Regression or Random Forests).
-For each run , we train from an independent random seed while keeping the data split, preprocessing, and hyperparameters fixed.
+For each run , we train under an independently instantiated training pipeline, while keeping the data split and preprocessing fixed.
+A pipeline instantiation may involve stochastic sources (e.g., data resampling or optimiser initialisation) and/or a sampled hyperparameter configuration, depending on the model class.
 After each training run, we compute feature-level attributions using a chosen interpretability method (here, SHAP), yielding an attribution vector over the input features.
 EvoXplain collects these attribution vectors,
-and treats them as samples drawn from the explanation distribution induced by stochastic training.
-Importantly, EvoXplain does not aggregate predictions or construct ensembles; it analyses explanation variability arising solely from repeated instantiation of the same model class.
+and treats them as samples drawn from the explanation distribution induced by repeated instantiation of the training pipeline.
+Importantly, EvoXplain does not aggregate predictions or construct ensembles; it analyses explanation variability arising solely from admissible variation within a single model class and data split.
+For deterministic solvers such as Logistic Regression with lbfgs, run-to-run variation arises from differences in admissible hyperparameter configurations rather than optimiser stochasticity.
+4.2 Orthogonal Synthetic Control Dataset
+To assess whether explanation multiplicity can arise purely from feature correlation or data redundancy, we introduce a synthetic control dataset with strictly orthogonal input features and a known linear signal structure.
+Features are generated to be mutually uncorrelated, and labels are assigned using a fixed ground-truth linear rule with additive noise.
+This setting represents a best-case scenario for explanation stability, where neither feature correlation nor redundant predictors can induce alternative attribution patterns.
+We apply the same repeated training and explanation extraction protocol used for real datasets.
+Under fixed regularisation strength, explanations collapse to a single explanatory regime ().
+When admissible training configurations are expanded to include a range of regularisation strengths, structured explanation multiplicity re-emerges despite strict feature orthogonality.
+This confirms that the observed multiplicity is not driven by feature correlation or attribution artefacts, but arises from non-identifiability in the admissible training pipeline itself.
 4.3 Representing Explanation Space
@@
 4.4 Clustering and Basin Structure
-To assess whether explanations converge to a single pattern or separate into multiple patterns, we apply -means clustering (chosen for simplicity and interpretability) for values of from to .
-For each , we compute the silhouette score and select
-If , explanations concentrate around a single explanatory mode. If , EvoXplain identifies multiple explanatory basins corresponding to distinct explanation patterns reached across training runs.
+To assess whether explanations converge to a single pattern or separate into multiple patterns, we evaluate candidate multi-basin solutions using -means clustering for values of .
+For each , we compute the silhouette score .
+We select
+subject to the constraint , where is a predefined silhouette threshold. If no candidate satisfies this criterion, the explanation distribution is treated as unimodal and we set .
+If this criterion is not met, explanations are treated as forming a single explanatory mode, even if under unconstrained maximisation.
+When the criterion is satisfied, EvoXplain identifies multiple regions of concentration in explanation space, which we refer to as explanatory basins for diagnostic convenience.
+We verified that conclusions are qualitatively stable for reasonable variations of (See Appendix A).
 4.5 Mechanistic Entropy and Basin Support
@@
 Clustering in explanation space is used here as a diagnostic tool rather than as a claim about the true number of underlying mechanisms.
-We adopt k-means with silhouette-based selection for simplicity and transparency, and because the resulting clusters are visually apparent in low-dimensional projections and consistent across independent train–test splits.
-Importantly, the presence or absence of explanation multiplicity does not hinge on fine-grained cluster boundaries: when training converges to a single explanatory regime (e.g., under fixed regularisation), explanations collapse to a single dense region and silhouette analysis consistently selects k = 1, whereas when multiple explanatory regimes are admissible, well-separated modes emerge with stable support across runs and splits.
+We adopt k-means with silhouette-based selection for simplicity and transparency, and because the resulting basin structure is visually apparent in low-dimensional projections and reproducible across independent train–test splits, even when the number or balance of basins varies.
+Importantly, the presence or absence of explanation multiplicity does not hinge on fine-grained basin boundaries: when training converges to a single explanatory regime (e.g., under fixed regularisation), explanations collapse to a single dense region and silhouette analysis consistently selects k = 1, whereas when multiple explanatory regimes are admissible, well-separated modes emerge with stable support across runs and splits.
 EvoXplain therefore does not rely on clustering to create structure, but uses clustering to quantify structure that is already present in the explanation distribution.
 4.9 Theoretical Interpretation
-EvoXplain can be understood through optimisation non-identifiability.
+EvoXplain can be understood through training-pipeline non-identifiability, which includes but is not limited to optimisation non-identifiability.
 Let denote the parameter space of a model class and its loss surface.
-In high-dimensional or weakly constrained settings, training may converge to multiple parameter configurations with near-optimal loss,
-which we refer to as a mechanism manifold. Different training runs may select different points in and yield distinct attribution patterns even when predictive performance is similar.
+In high-dimensional or weakly constrained settings, or when the training pipeline admits multiple admissible configurations, training may converge to multiple parameter configurations with near-optimal loss,
+which we refer to as a mechanism manifold. Different training instantiations may select different points in and yield distinct attribution patterns even when predictive performance is similar.
 Let map model parameters to attribution vectors. EvoXplain studies the empirical distribution
@@
 When , training concentrates near a single attractor; when , training repeatedly reaches different regions of the explanation manifold.
-EvoXplain does not claim these basins correspond to true causal mechanisms. Rather, it demonstrates that optimisation and model capacity alone can induce multiple, stable explanatory behaviours.
+EvoXplain does not claim these basins correspond to true causal mechanisms.
+Rather, it demonstrates that non-identifiability in the training pipeline—including optimisation dynamics, model capacity, and admissible configuration choices—can induce multiple, stable explanatory behaviours.
 4.10 Computational Considerations
@@
 5 Scope and Limitations
-EvoXplain does not assert that the clusters it identifies correspond to true causal mechanisms, nor that particular centroids represent any form of ground truth.
+EvoXplain does not assert that the basins it identifies correspond to true causal mechanisms, nor that particular centroids represent any form of ground truth.
 Its outputs are contingent upon the attribution method used, and the observed structure may also reflect noise or inductive biases inherent to the model or the explanation technique.
@@
 EvoXplain has been designed to assess how models behave across repeated training runs.
-It determines whether a model settles into a single cluster of explanations or whether different explanations emerge across retraining.
+It determines whether a model settles into a single basin of explanations or whether different explanations emerge across retraining.
 This perspective comes with several limitations.
@@
 For large-scale models, such as foundation systems, full retraining may be impractical, and alternative strategies—such as checkpoint sampling or multiple fine-tuning runs—may be required.
+SHAP is used here as a widely adopted probe of explanatory structure rather than as a privileged or canonical explanation method; EvoXplain evaluates the stability of attribution geometry under retraining, not the correctness of any particular attribution paradigm.
 These limitations do not diminish the framework, but rather clarify the scope of its current applicability.
@@
 6.3 Repeated Training Protocol
-For each of the 10 dataset–model–split combinations (with five splits used for Random Forests), we perform independent training runs.
+For each dataset–model–split combination (10 splits for Logistic Regression and 5 splits used for Random Forests), we perform training runs.
 Each run differs only in:
@@
 the regularisation or model hyperparameters when explicitly varied.
+For Logistic Regression with varied , repeated runs correspond to repeated grid assignments (not stochastic optimisation), so variability reflects admissible pipeline variation along the regularisation path rather than run-to-run optimiser randomness.
 For Logistic Regression, the regularisation strength is either:
 fixed (), or
-varied deterministically over a log-spaced grid between and .
+varied deterministically by assigning each run index to a value in a fixed log-spaced grid , cycling through as increases (so grid values repeat across the runs).
 For Random Forests, hyperparameters are sampled from predefined grids over: number of estimators, maximum depth, minimum samples per leaf, and maximum features.
 Each training run produces a trained model, test accuracy, and a global explanation vector.
-Runs are executed in parallel on HPC infrastructure using chunked execution and aggregated post hoc (See the code repository).
+Runs are executed in parallel on HPC infrastructure using chunked execution and aggregated post hoc (exploratory analyses in the repository).
 Due to the computational cost of Random Forest explanations, RF experiments were conducted over five independent train–test splits, which were sufficient to characterise the qualitative geometry of the explanation landscape.
-Although the lbfgs solver is deterministic for fixed data and fixed hyperparameters, explanation multiplicity arises because training spans a range of regularisation strengths, allowing optimisation to settle into different equally optimal basins across the hyperparameter-conditioned loss landscape.
-Thus, the stochasticity measured by EvoXplain reflects basin selection under admissible training configurations, not randomness in optimisation or attribution.
+Although the lbfgs solver is deterministic for fixed data and fixed hyperparameters, explanation multiplicity arises because the training pipeline spans a range of admissible regularisation strengths, each inducing a distinct optimal solution along the regularisation path.
+Thus, the variability measured by EvoXplain reflects selection among equally admissible solutions under the specified training configuration, rather than randomness in optimisation or attribution.
 6.4 Explanation Generation
@@
 Explanation vectors are -normalised prior to clustering. Degenerate cases (nearly identical vectors) are detected using a fixed tolerance.
-Clustering is performed using -means with candidate values
-The optimal number of clusters is selected by maximising the silhouette score. Cluster centroids, sizes, and mechanistic entropy are computed as described in 4.3.
+We treat a single basin () as the null hypothesis corresponding to a unimodal explanation distribution.
+For each candidate , we compute the silhouette score .
+A multi-basin solution is accepted only if , where we set as a conservative guardrail against over-partitioning.
+If no candidate meets this criterion, the explanation distribution is treated as unimodal and we set .
+In all experiments reported here, the minimum silhouette score among accepted multi-basin solutions exceeded 0.38, so conclusions do not hinge on the specific value of .
+To assess robustness to clustering choices, we additionally replicate basin detection using alternative clustering metrics and linkage criteria (Appendix A).
+When one or more admissible multi-basin solutions exist, the optimal number of basins is selected by maximising the silhouette score among admissible .
+Basin centroids, sizes, and mechanistic entropy are then computed as described in Section 4.4.
+While k-means assumes approximately spherical clusters, we use it here as a pragmatic heuristic for detecting multimodality in explanation space rather than as a geometric model of the manifold itself.
+In all reported cases, accepted multi-basin solutions exhibit high silhouette scores, indicating well-separated modes despite manifold curvature.
 6.6 Disagreement Inspection
-When , we identify representative runs for each explanatory basin by selecting the run closest to the cluster centroid in cosine distance.
-Representative models are retrained and evaluated on the same test set. For each test instance , we compute the disagreement score
-where denotes a representative model from basin .
+When , we identify representative runs for each explanatory basin by selecting the run closest to the basin centroid in cosine distance.
+The corresponding representative trained models are evaluated on the same test set. For each test instance , we compute the disagreement score
+where denotes a representative model instance drawn from basin .
 Instances with the largest are reported to characterise cases where distinct explanatory modes lead to materially different predictions.
 6.7 Visualisation
-For each split, explanation vectors are projected into two dimensions using PCA after centring. Per-split manifolds display cluster structure and regularisation gradients.
+For each split, explanation vectors are projected into two dimensions using PCA after centring. Per-split manifolds display basin structure and regularisation gradients.
 A universal manifold is constructed by stacking explanation vectors across all splits and applying PCA without per-split centring, enabling comparison of global geometric structure across data realisations.
 7 Results
-7.1 Logistic Regression: explanation multiplicity under stochastic training
+7.1 Logistic Regression: explanation multiplicity under admissible pipeline variation
 7.1.1 Breast Cancer
@@
 When regularisation strength is allowed to vary, this expectation does not hold.
-For a representative split (split seed 109), we analyse 1,000 independent training runs.
+For a representative split (split seed 109), we analyse 1,000 runs under the repeated training protocol.
 In explanation space, the SHAP vectors do not form a single cloud.
-Instead, they organise along a smooth, low-dimensional arc with clear separation between regions (Fig.
+Instead, for Logistic Regression they organise along a smooth, low-dimensional explanation manifold that traces a curved arc, within which recurrent regions of high density emerge with clear separation between them (Fig.
 1).
@@
 The normalised explanation entropy is 0.922, showing that runs are distributed across the two modes rather than collapsing into a single dominant explanation.
-The two clusters are not small perturbations of one another.
-Cluster 0 contains 663 runs with mean , while Cluster 1 contains 337 runs with mean .
-Despite occupying different regions of the explanation manifold, both clusters achieve comparable predictive accuracy (approximately 0.96).
-The accuracy distributions overlap substantially (Fig.
-1, right panel).
+The two basins are not small perturbations of one another.
+Basin 0 contains 663 runs with mean , while Basin 1 contains 337 runs with mean .
+Despite occupying different regions of the explanation manifold, both basins achieve comparable predictive accuracy (approximately 0.96, See Table 1).
 In other words, both explanation modes achieve similarly high predictive performance.
-The clusters correspond to distinct regularisation regimes.
-The right panel of Fig.
-1 shows a smooth gradient of values along the arc: strong regularisation (low ) concentrates at one end, while weak regularisation (high ) occupies the other.
-This gradient structure indicates that explanation geometry is tightly coupled to the regularisation–Loss landscape interaction.
+Although regularisation strength varies smoothly along the explanation manifold (Fig.
+1, right), the two explanation basins are not simple partitions of the sweep.
+Both basins span the full sampled range of regularisation values (Table 2), and runs with very similar can occupy different regions of the manifold.
+This overlap persists even when conditioning on narrow bands of regularisation strength, indicating that regularisation correlates with the continuous geometry of explanation space but does not uniquely determine basin membership.
+The mechanistic multiplicity therefore reflects non-identifiability under admissible training configurations, rather than trivial hyperparameter binning or noise.
 These differences are reflected in the explanations themselves.
-The dominant features differ qualitatively between clusters.
-One mode emphasises classical size-related measurements, while the other places substantially more weight on variability- and shape-related features.
-The model is not merely reweighting the same features slightly differently; it is relying on qualitatively different explanatory patterns.
-Importantly, the average attribution vector—computed as the arithmetic mean of all SHAP vectors across runs—does not lie on the populated manifold.
-In the split-level embedding, it falls in a sparsely populated region between clusters (Fig.
-1), indicating that this average explanation does not correspond to a typical trained model instance.
-When explanation vectors are stacked across all ten splits, the same qualitative picture persists.
-The universal explanation manifold shows ten parallel arcs, each corresponding to a different train–test realisation, that remain separated in explanation space, while predictive accuracy remains uniformly high along all tracks (Fig.
-2).
-The global average attribution vector, computed across all runs and all splits, lies again away from the high-density regions of the manifold.
-This reinforces that averaging explanations across stochastic runs produces a vector that is not representative of any deployed model instance.
-To assess whether these explanation differences translate into different predictions at the individual level, we compare representative models from the two clusters in split 109.
-In all held test instances, the mean absolute disagreement in the predicted malignant probability is 0.067 and the median is 0.035.
-However, the tail of the distribution is substantial: the maximum disagreement reaches , with 34 instances exceeding , 12 instances exceeding , and 4 instances exceeding .
-For the highest-disagreement case, one representative model assigns a malignant probability substantially different from the other, despite both models achieving high overall accuracy.
-These discrepancies arise even though aggregate performance metrics are indistinguishable.
+The dominant features differ across basins: one mode places greater emphasis on classical size-related measurements, while the other assigns relatively more weight to variability and shape-related features.
+While these patterns arise within the same feature space, the resulting attribution profiles exhibit distinct explanatory emphases rather than simple uniform rescaling of a single feature ranking.
+Importantly, the average attribution vector—computed as the arithmetic mean of all SHAP vectors across runs—lies in a low-density region.
+In the split-level embedding, it falls in a sparsely populated region between basins (Fig.
+2), indicating that this average explanation does not correspond to any high-density region of the explanation distribution.
+When explanation vectors are stacked across all ten train–test splits and visualised jointly (without performing clustering), the same qualitative picture persists.
+In the resulting embedding, each split traces a distinct arc in explanation space corresponding to its specific train–test realisation (Fig.
+1).
+These arcs remain separated from one another, even though predictive accuracy remains uniformly high along all trajectories.
+To assess whether these explanation differences translate into different predictions at the individual level, we compare representative models from the two basins in split 107.
+While the mean absolute disagreement in predicted malignant probability across test instances is modest, the upper tail of the distribution is substantial.
+In this split, 76 test instances exhibit disagreement exceeding , 33 exceed , and 15 exceed , with the maximum observed disagreement reaching .
+For the highest-disagreement case, one representative model assigns a malignant probability of 0.72 while the other assigns 0.04, leading to opposite predicted labels under a 0.5 threshold, despite both models achieving similarly high overall accuracy.
+These discrepancies arise even though aggregate performance metrics are effectively indistinguishable.
 Control Experiment: Stability under Fixed Regularisation.
 As a control, we repeat the same analysis under fixed regularisation strength ().
-Across ten independent train–test splits, each with 1,000 runs, the explanation vectors consistently collapse to a single cluster (), with negligible dispersion in explanation space.
-This confirms that the observed explanation multiplicity is not an artefact of the SHAP method or dimensionality reduction, but emerges specifically when the training process admits multiple equally accurate models.
+Across ten independent train–test splits, each with 1,000 runs, the explanation vectors consistently collapse to a single basin (), with negligible dispersion in explanation space.
+This indicates that, in this setting, explanation multiplicity arises under admissible hyperparameter-conditioned training configurations (varied ), rather than from randomness in the optimiser or attribution method.
+(This is true for Adult Income dataset as well, which can be found in Appendix A)
+Orthogonal Synthetic Control.
+To rule out feature correlation as a driver of explanation multiplicity, we repeat the analysis on a synthetic dataset with strictly orthogonal features and a known linear ground-truth rule.
+Under fixed regularisation, explanations collapse to a single mode ().
+When regularisation is allowed to vary across admissible values, explanations again organise into a structured low-dimensional manifold with recurrent regions, despite the absence of correlated or redundant features.
+This demonstrates that explanation multiplicity does not rely on feature correlation or dataset-specific quirks, but can arise even in idealised settings where explanatory ambiguity would not ordinarily be expected.
+Full details are provided in Appendix A.
 | Split | Runs | Silh. () | Acc. mean | Acc. std | ||
@@
 | 109 | 1 | 337 | 12.199 | 21.267 | [0.011, 98.3] | 0.962 | 0.007 |
-Across splits, Logistic Regression on the Breast Cancer dataset consistently exhibits two well-separated explanation basins under stochastic training (Table 1).
-In every split, clustering selects explanation modes with high silhouette scores (0.76–0.82) and near-maximal normalised entropy (0.91–0.94), indicating that runs are distributed across multiple explanatory mechanisms rather than collapsing into a single dominant solution.
+A potential alternative explanation is that the observed basins merely reflect different regularisation regimes.
+However, this is not supported by the data.
+As shown in Table 2, the two explanation basins identified for Split 100 have nearly identical mean regularisation strengths (), yet occupy well-separated regions of explanation space.
+This demonstrates that explanation multiplicity is not reducible to a monotonic effect of regularisation strength, but instead reflects distinct explanation modes compatible with similar model configurations.
+Across splits, Logistic Regression on the Breast Cancer dataset consistently exhibits two well-separated explanation basins under admissible training pipeline variation (Table 1).
+In every split, clustering selects explanation modes with high silhouette scores (0.76–0.82) and near-maximal normalised entropy (0.91–0.94), indicating that runs are distributed across multiple explanation basins rather than collapsing into a single dominant pattern.
 This structure is stable across independent train–test realisations.
-Importantly, predictive performance remains uniformly high across splits and clusters.
+Importantly, predictive performance remains uniformly high across splits and basins.
 Mean accuracy ranges from 0.956 to 0.976, with small standard deviations, and accuracy distributions overlap substantially between explanation basins within each split.
 There is no split in which explanation multiplicity coincides with degraded or unstable predictive performance, consistent with the split-level and universal explanation manifolds.
-Cluster-level analysis (Table 2) further shows that the two explanation basins are not reducible to a simple partition by regularisation strength.
-Within each split, both clusters span the full sampled range of values, with similar mean and variance.
+Basin-level analysis (Table 2) further shows that the two explanation basins are not reducible to a simple partition by regularisation strength.
+Within each split, both basins span the full sampled range of values, with similar mean and variance.
 This indicates that the mechanistic separation observed in explanation space reflects genuinely distinct explanatory modes rather than trivial hyperparameter binning.
-Note that while position along the arc correlates smoothly with regularisation strength (Fig.1, right panel), cluster membership does not partition by C: runs with identical regularisation can land in different basins depending on other aspects of the optimisation trajectory.
-As a control, when regularisation strength is fixed, explanations collapse to a single cluster across all splits, confirming that explanation multiplicity emerges specifically when the training process admits multiple equally accurate solutions.
+As a control, when regularisation strength is fixed, explanations collapse to a single basin across all splits, confirming that explanation multiplicity emerges specifically when the training process admits multiple equally accurate solutions.
 7.1.2 COMPAS
 Results for Logistic Regression on the COMPAS dataset are summarised in Figures 3 and 4 and in Tables 3 and 4.
-Across all ten train–test splits, explanation clustering consistently selects two clusters (), with silhouette scores ranging from 0.73 to 0.91 (Table 3).
+Across all ten train–test splits, explanation clustering consistently selects two basins (), with silhouette scores ranging from 0.73 to 0.91 (Table 3).
 These values indicate strong and stable separation between explanation modes across resampled datasets—notably higher than those observed on Breast Cancer.
-The normalised entropy ranges from 0.83 to 0.91, indicating that explanation modes are somewhat unequally populated, with cluster sizes typically in a ratio of approximately 1:2.5.
+The normalised entropy ranges from 0.83 to 0.91, indicating that explanation modes are somewhat unequally populated, with basin sizes typically in a ratio of approximately 1:2.5.
 Predictive performance remains stable across splits.
-Mean accuracy ranges from 0.68 to 0.70 (this is typical for recidivism prediction tasks given inherent unpredictability), with standard deviations on the order of or smaller (Table 3).
+Mean accuracy ranges from 0.68 to 0.70 (consistent with typical reported performance ranges for this task), with standard deviations on the order of or smaller (Table 3).
 There is no evidence that the emergence of multiple explanation modes is associated with unstable or degraded performance.
 This is consistent with the split-level manifolds (Fig.
-3), where accuracy varies smoothly along the explanation geometry rather than separating sharply by cluster.
-Table 4 shows that, as with Breast Cancer, both explanation clusters span the full range of regularisation strengths.
-In every split, both clusters contain runs with values ranging from 0.01 to approximately 100, with similar mean values of around 10–12.
+3), where accuracy remains confined to a narrow range across the explanation geometry rather than separating sharply by basin.
+Table 4 shows that, as with Breast Cancer, both explanation basins span the full range of regularisation strengths.
+In every split, both basins contain runs with values ranging from 0.01 to approximately 100, with similar mean values of around 10–12.
 This indicates that the mechanistic separation in explanation space is not a simple partition by regularisation strength.
 Rather, the two explanation modes represent genuinely distinct regions of the explanation landscape that models can reach regardless of their regularisation setting.
-Predictive accuracy remains effectively identical across clusters within each split, with differences typically below 0.001.
+Predictive accuracy remains effectively identical across basins within each split, with differences typically below 0.001.
 When explanation vectors are stacked across all splits, the same structure persists (Fig.
@@
 As a control, we repeat the same analysis under fixed regularisation strength ().
-Across all ten independent train–test splits, each with 1,000 runs, the explanation vectors consistently collapse to a single cluster (), with zero entropy and negligible dispersion in explanation space.
+Across all ten independent train–test splits, each with 1,000 runs, the explanation vectors consistently collapse to a single basin (), with zero entropy and negligible dispersion in explanation space.
 This confirms that the observed explanation multiplicity is not an artefact of the SHAP method or clustering procedure, but emerges specifically when the training process admits multiple equally accurate models through varied regularisation.
@@
 To assess whether explanation multiplicity can translate into materially different predictions at the individual level, we compare representative models drawn from distinct explanation basins within the same train–test split.
-For each held-out test instance , we compute the absolute difference in predicted malignant probability,
-where and denote representative models selected from two different explanation clusters.
+For each test instance , we compute the disagreement score
+where denotes a representative model selected from explanatory basin . In the present experiments, clustering yields basins, so this reduces to
 Importantly, the compared models share the same model class, training data, and optimisation objective, and achieve comparable test accuracy.
-This controls for performance differences and isolates disagreement attributable to distinct explanatory mechanisms rather than optimisation noise or underfitting.
+This controls for trivial performance differences, ensuring that observed disagreement reflects differences in admissible explanatory behaviour under the training pipeline rather than underfitting or attribution noise.
 On the Breast Cancer dataset, disagreement magnitudes span a wide range.
@@
 These values indicate that explanation multiplicity can manifest as large probability divergences for individual patients even when aggregate performance metrics are indistinguishable.
-The highest-disagreement case in Figure 5 corresponds to a malignant tumour for which one model assigns a probability of 0.72 while the other assigns 0.04, leading to opposite clinical decisions for the same patient.
+Split 107 is shown as an illustrative high-disagreement example; other splits exhibit similar behaviour with varying magnitude.
+The highest-disagreement case in Figure 5 corresponds to a malignant tumour for which one model assigns a probability of 0.72 while the other assigns 0.04, leading to opposite predicted labels under a 0.5 threshold for the same patient.
 Notably, this disagreement does not arise from marginal predictions near the decision threshold: one model is confidently malignant, while the other is confidently benign.
@@
 Although both models attend to overlapping morphological feature families, they weight and combine this evidence differently, resulting in incompatible diagnostic conclusions.
-This behaviour is inconsistent with an interpretation of disagreement as stochastic fluctuation or local instability, and instead reflects the presence of distinct explanatory mechanisms that remain viable at comparable accuracy.
+This pattern persists across multiple representative pairs within the same basins (exploratory analyses in the repository).
 Across all ten train–test splits, the magnitude of individual-level disagreement varies substantially.
@@
 This variability indicates that the practical manifestation of explanation multiplicity depends on the particular data realisation.
-However, critically, large individual-level disagreements can arise even when aggregate accuracy and calibration metrics remain effectively unchanged.
+However, critically, large individual-level disagreements can arise even when aggregate accuracy and aggregate performance metrics remain effectively unchanged.
 In such cases, standard performance summaries obscure the fact that different model instances may assign substantially different risk estimates to the same individual.
 Crucially, disagreement is not confined to cases where models predict different labels.
-Table 5 reports representative test instances from Split 107 for which both explanation basins correctly predict the same class, yet assign substantially different malignant probabilities ().
+Table 5 reports representative test instances from Split 107 for which both explanation basins correctly predict the same class, yet assign substantially different malignant probabilities (–).
 In these cases, predictions agree at the label level, but differ meaningfully in confidence, with one model indicating borderline risk and the other indicating high certainty.
-While these instances do not lead to immediate label flips, they demonstrate that mechanistic disagreement can persist even when classification outcomes are identical, manifesting as divergent belief states rather than discrete decision changes.
+While these instances do not lead to immediate label flips, they demonstrate that explanation-level disagreement can persist even when classification outcomes are identical, manifesting as divergent confidence assessments rather than discrete decision changes.
 Such differences are relevant in downstream clinical contexts where predicted probabilities inform risk stratification, follow-up decisions, or human oversight.
@@
 When multiple explanation basins coexist at similar accuracy, the choice of model instance can materially affect individual-level predictions, either by altering the predicted class or by substantially shifting estimated risk.
-The observed clusters therefore correspond to qualitatively distinct explanatory mechanisms, not small perturbations of a single underlying solution.
+The observed basins therefore correspond to qualitatively distinct explanatory mechanisms, not small perturbations of a single underlying solution.
 High predictive performance alone does not imply mechanistic stability or explanatory uniqueness.
@@
 | Test index | True label | |||
-| 36 | Benign | 0.79 | 1.00 | 0.21 |
-| 169 | Malignant | 0.60 | 0.81 | 0.21 |
+| 29 | Malignant | 0.85 | 0.60 | 0.25 |
+| 136 | Benign | 0.23 | 0.00 | 0.23 |
 7.3 Random Forests: explanation multiplicity in ensemble models
-Having established that explanation multiplicity arises in Logistic Regression under stochastic training—despite its convex objective and stable predictive performance—we next examine whether the same phenomenon appears in a qualitatively different model class.
-Random Forests introduce non-convexity, ensemble averaging, and additional sources of stochasticity through bootstrap sampling and feature subsampling.
-Evaluating explanation stability in this setting allows us to test whether the observed multiplicity is specific to linear models under hyperparameter variation, or whether it persists in more expressive and widely deployed ensemble architectures.
+We next evaluate EvoXplain on Random Forests, a model class that introduces nonlinearity, ensemble averaging, and intrinsic stochasticity through bootstrap resampling and feature subsampling.
+This setting allows us to test whether explanation multiplicity persists when predictive variance is strongly suppressed by ensembling, using five independent train–test splits.
 7.3.1 Breast Cancer
@@
 7.
-Each train–test split aggregates approximately Random Forest instances, reflecting stochasticity arising from bootstrap resampling, feature subsampling, and hyperparameter variation.
-This large-scale sampling allows the structure of explanation space to be examined at high resolution.
+We train a population of Random Forest instances per train–test split by sampling stochastic sources intrinsic to the algorithm (bootstrap resampling and feature subsampling).
 Figure 6 visualises the universal explanation manifold obtained by stacking explanation vectors across multiple train–test splits.
-In contrast to the smooth, curved manifolds observed for Logistic Regression, Random Forest explanation vectors organise into multiple dense and well-separated regions.
+In contrast to the smooth, curved manifolds observed for Logistic Regression, Random Forest explanation vectors organise into two well-separated regions in explanation space, with partial overlap at their boundaries.
 These regions persist across splits, indicating stable explanation multiplicity in the ensemble setting.
-As in the linear case, the global average attribution vector—computed by averaging SHAP values across all runs—lies outside the high-density regions of the manifold.
-This indicates that aggregating explanations across stochastic training runs again produces a representation that does not correspond to a typical deployed model instance.
-Importantly, predictive accuracy varies smoothly across the explanation manifold and does not align with any single cluster or region.
-Models drawn from different explanation clusters achieve comparable test accuracy, with no sharp performance boundary separating the modes.
+As in the linear case, the global average attribution vector—computed by averaging SHAP values across all instances—lies outside the high-density regions of the manifold.
+This indicates that aggregating explanations across admissible training pipeline instantiations produces a representation that does not correspond to any high-density region of the explanation distribution.
+Importantly, predictive accuracy does not align with any single basin or region.
+Instances drawn from different explanation basins achieve comparable test accuracy, with no sharp performance boundary separating the modes.
 This demonstrates that, in Random Forests, explanation multiplicity is not driven by accuracy trade-offs or poorly performing solutions, but instead reflects the coexistence of distinct explanatory mechanisms within the same performance regime.
 Figure 7 shows the explanation landscape for a single representative split (Split 104).
-Even within a fixed train–test realisation, explanation vectors again separate into two high-density regions.
-The first principal component captures the majority of variance in explanation space, while the second component reflects residual spread within each mode.
-As in the universal manifold, predictive accuracy varies continuously across both regions, indicating that explanation multiplicity arises intrinsically within a single split and does not depend on aggregation across data resamplings or extreme hyperparameter configurations.
+Even within a fixed train–test realisation, explanation vectors again form two statistically separable regions in the explanation space.
+The first principal component captures the majority of the variance in explanation space, while the second component reflects the residual spread within each mode.
+As in the universal manifold, predictive accuracy remains comparable across these regions, indicating that explanation multiplicity is already present within a single train–test split and is not an artefact of aggregating explanations across multiple data resamplings.
 Taken together, these results show that explanation multiplicity persists in Random Forests despite ensemble averaging and strong inductive bias towards prediction stability.
@@
 This suggests that interpretability cannot be inferred from model form or accuracy alone.
-For example, linear models commonly assumed to be transparent do not consistently converge to a single explanatory mechanism under standard retraining procedures.
-In our experiments, Logistic Regression exhibits explanation multiplicity across datasets when regularisation is allowed to vary, despite high and stable predictive accuracy (Lipton (2018); Rudin (2019)).
+For example, linear models commonly assumed to be transparent do not consistently converge to a single explanatory mechanism under standard retraining and model selection procedures.
+In our experiments, Logistic Regression exhibits explanation multiplicity across datasets when the training pipeline admits multiple admissible configurations (here instantiated via regularisation variation), despite high and stable predictive accuracy (Lipton, 2018; Rudin, 2019).
 This challenges the common practice of treating transparency and accuracy as sufficient conditions for interpretability.
@@
 While predictive performance remains stable, explanation multiplicity matters because explanations are often used to justify, audit, or regulate model-driven decisions at the individual level.
-This suggests that, rather than just focusing on a single model instance, we need to develop stability-aware interpretability of models that provides explanations based on the behaviour of models over multiple instances or runs (Arrieta and others (2020); Selbst and Barocas (2018)).
-EvoXplain does not claim to identify where multiplicity originates—whether from data structure, model flexibility, or optimisation randomness.
+This suggests that, rather than just focusing on a single model instance, we need to develop stability-aware interpretability of models that provides explanations based on the behaviour of models over multiple instances or runs (Arrieta and others, 2020; Selbst and Barocas, 2018).
+EvoXplain does not claim to identify where multiplicity originates—whether from data structure, model flexibility, or the non-identifiability of admissible training configurations within a pipeline.
 Instead, it makes this uncertainty visible.
 The framework is deliberately agnostic to causal origin and instead focuses on empirical detectability.
+Deterministic models can still exhibit pipeline-induced multiplicity.
+A potential source of confusion is that regularised Logistic Regression with a deterministic solver is uniquely determined by a fixed dataset, preprocessing, and a fixed regularisation strength .
+EvoXplain does not assume that a deployed system samples only from this single point.
+In practice, retraining and model updates include upstream degrees of freedom such as hyperparameter selection, data refresh, and preprocessing changes, which together induce a distribution over admissible model instantiations (D’Amour et al., 2022; Breiman, 2001; Marx et al., 2020).
+Consistent with this framing, an orthogonal synthetic control confirms that with fixed explanations collapse to , whereas allowing to vary reintroduces multimodality, indicating that multiplicity arises from the instantiation pipeline rather than solver randomness.
+Logistic Regression is the minimal test case.
+It is important to emphasise that -regularised Logistic Regression represents the best-case scenario for explanation stability: the loss is convex, the solver is deterministic, and the model is fully identifiable once the regularisation strength is fixed.
+If explanation multiplicity arises even in this setting when the deployment pipeline admits multiple admissible instantiations, then the phenomenon is necessarily more pronounced in model classes with intrinsic stochasticity or non-convexity.
+Random Forests introduce bootstrap and feature-subsampling randomness; Elastic Net regularisation admits multiple sparse solutions under correlated features; and deep neural networks exhibit vast equivalence classes of parameterisations with similar predictive behaviour.
+EvoXplain therefore establishes a lower bound on explanation multiplicity: if a pipeline fails to identify a unique explanatory regime in the simplest convex case, it cannot be expected to do so in more flexible or stochastic architectures.
+Importantly, the presence of multiple explanation basins at comparable hyperparameter values—for example, basins with nearly identical mean regularisation strength in Logistic Regression (Table 1)—rules out an interpretation of EvoXplain as merely measuring continuous attribution drift, and instead points to discrete explanation non-identifiability within admissible training pipelines.
 The multiplicity observed cannot be attributed to instability or noise in SHAP itself.
 SHAP values are deterministic for any fixed model.
-EvoXplain examines attribution profiles across independent model trainings, rather than variability within a single explanation.
+EvoXplain examines attribution profiles across independently trained model instances produced by the same pipeline, rather than variability within a single explanation.
 This distinction is critical: EvoXplain characterises variability across trained models, not variability within an attribution method.
-Because SHAP is deterministic for a fixed model, any observed clustering reflects differences between trained models rather than stochasticity in the explainer itself.
-As such, the resulting clusters reflect real differences in learned mechanisms.
+Because SHAP is deterministic for a fixed model, any observed clustering reflects differences between trained model instances rather than stochasticity in the explainer itself.
+As such, the resulting basins reflect real differences in learned parameterisations and their induced attribution behaviour.
 For the Breast Cancer benchmark, the centroids of the discovered basins reinforce established diagnostic attributes such as perimeter, radius, area, concavity, and associated morphological descriptors used in clinical tumour assessment.
@@
 In this sense, explanation multiplicity is a property to be measured, reported, and reasoned about, rather than suppressed.
-In the presence of multiple modes of explanation, the decision to select one becomes a question of science or governance, rather than something the training pipeline can resolve automatically (European Commission (2023); Rudin (2019)).
-This perspective aligns with recent safety assurance methodologies for machine learning, such as AMLAS and the “BIG Argument” for AI safety proposed by Habli and colleagues (Hawkins et al.
-(2021); Habli et al.
-(2025)), which emphasise the need for explicit evidence of model robustness beyond predictive accuracy alone.
+In the presence of multiple modes of explanation, the decision to select one becomes a question of science or governance, rather than something the training pipeline can resolve automatically (European Commission, 2023; Rudin, 2019).
+This perspective aligns with recent safety assurance methodologies for machine learning, such as AMLAS and the “BIG Argument” for AI safety proposed by Habli and colleagues (Hawkins et al., 2021; Habli et al., 2025), which emphasise the need for explicit evidence of model robustness beyond predictive accuracy alone.
 Within such frameworks, explanation stability is not merely an interpretability preference but a form of verification evidence.
-EvoXplain contributes precisely this type of evidence by quantifying how model reasoning varies under standard retraining procedures, thereby supporting or challenging claims of reliability in safety-critical ML systems (Paterson et al.
-(2025)).
+EvoXplain contributes precisely this type of evidence by quantifying how model reasoning varies under standard retraining and model selection procedures, thereby supporting or challenging claims of reliability in safety-critical ML systems (Paterson et al., 2025).
 A closely related operational response to model multiplicity is to aggregate explanations over a set of near-optimal models and report uncertainty around the aggregate.
@@
 In such settings, interpretability requires representing the set of mechanistic regimes and their relative support (frequency), rather than collapsing them into a single averaged explanation with uncertainty bands.
-This is particularly important because explanations have immediate consequences in certain domains.
-In criminal justice contexts, multiple basins of attraction counteract the assumption that a model provides a singular, stable justification Umbach et al.
-(2023).
-In biomedicine, instability observed even on standard datasets warns that attribution-based interpretations should be treated with caution when used to support clinical claims Agrawal and Jain (2023).
-Thus, EvoXplain helps to identify cases where a single explanation is insufficient to account for the behaviour of a model under retraining.
-In both settings, explanation stability becomes relevant precisely because explanations are used as part of decision-making processes rather than as post-hoc descriptions.
-In our experiments, we focus on tabular learning, where repeated training is manageable.
-However, retraining large language models (LLMs) presents a more difficult challenge.
-We do not claim that the specific geometric structures observed here generalise directly to large-scale models; rather, we argue that the underlying issue of mechanistic non-identifiability is likely to be amplified in such settings.
-LLMs train in very high-dimensional, non-convex spaces, are exposed to stochastic data, and follow fine-tuning paths that may diverge substantially.
-Whether models that exhibit similar external behaviour in such settings rely on the same internal mechanisms remains an open empirical question.
-Measuring this at scale is challenging, but not impossible: models and their fine-tuned variants can be compared, checkpoints sampled, and attribution patterns analysed across independent runs (Olah et al.
-(2020)).
-EvoXplain provides a conceptual starting point.
-The instability observed even in small models suggests that non-identifiability in foundation models may be important and deserves further study (Lakshminarayanan et al.
-(2017)).
-Beyond the empirical evidence, mechanistic multiplicity underscores the need to evaluate reliability and interpretability as properties that cannot be assumed.
-Instead, they must be empirically established under realistic training variability.
-Why Mechanistic Multiplicity Matters
-Mechanistic multiplicity primarily raises questions of reliability, rather than philosophy.
-It concerns whether a modelling pipeline produces consistent reasoning under retraining, not whether a single explanation appears plausible.
-When predictive accuracy decays, retraining is standard practice, and the modelling process may generate different reasoning outputs, with no guarantee of stable reasoning over time.
-Reliability must therefore concern the process rather than a particular model instance.
-In high-stakes contexts, explanations justify actions.
-Even at equal predictive accuracy, models can diverge in the actions, risk factors, and scientific hypotheses they imply, settling into different basins of explanation.
-This divergence is not resolved by appeals to sparsity, simplicity, or accuracy alone.
-When explanations inform decisions rather than merely describe them, instability in the underlying mechanism becomes a substantive safety and interpretability issue.
-Appeals to simplicity do not resolve multiplicity.
-In the presence of competing explanation basins with equivalent sparsity, complexity, and predictive accuracy, Occam’s razor cannot favour one over another.
-What remains is a form of empirical underdetermination.
-In many cases, competing explanation basins are comparable not only in accuracy, but also in sparsity and apparent simplicity, leaving no principled criterion for selection based solely on model-internal properties.
-In such cases, the presence of multiple stable explanatory modes evidences a profound underdetermination, which EvoXplain is designed to identify and document.
-There are boundaries to this work.
-We analyse explanation stability using attribution methods; other interpretability approaches, such as causal abstraction or circuit tracing, may offer different perspectives (Olah et al.
-(2020)).
-EvoXplain does not attempt to correct bias, determine which basin is “correct”, or quantify downstream risks such as fairness or failure modes Selbst and Barocas (2018).
-We consider only a limited number of datasets and model classes; this study represents an initial step toward broader analysis across architectures, attribution schemes, and domains such as reinforcement learning or language modelling.
-These limitations define the scope of the present study rather than its validity.
-The contribution of EvoXplain is not to propose a new explainer, but to provide a diagnostic framework that makes explanation instability measurable, comparable, and reportable across model populations.
-Despite these limitations, EvoXplain opens new research questions.
-Are explanation basins correlated with latent subgroups, causal mechanisms, or inductive bias pathways (Marx et al.
-(2020))?
-Is it possible to prefer certain basins for safety, fairness, or generalisability (Rudin (2019))?
-More generally, EvoXplain shifts interpretability from analysing individual models to analysing populations of models trained under the same procedure.
-This reframing allows explanation stability to be measured, compared, and monitored at the level of model populations.
-To summarise, EvoXplain treats interpretability as a property of model behaviour under retraining.
-By focusing on mechanistic multiplicity rather than assuming a single explanatory mode, EvoXplain provides tools to uncover and characterise multiplicity when it arises.
-This, in turn, raises methodological, epistemic, and governance challenges.
-If models admit multiple internally consistent mechanistic explanations, then explanation selection, validation, and governance require explicit criteria.
-If models embody a multiplicity of mechanistic stories, how should they be assessed, described, or governed?
-While EvoXplain does not resolve these questions, it makes them explicit and positions explanation stability as a concrete target for future research.
-In practice, EvoXplain enables practitioners and auditors to detect when single-model explanations are insufficient, without prescribing which explanation should be preferred.
 9 Conclusion
 Explanations are often treated as fixed model properties—examined once and assumed to reflect one underlying mechanism.
-EvoXplain argues the opposite, and instead quantifies this assumed stability.
+EvoXplain argues the opposite, and instead quantifies this assumed stability under realistic retraining and model selection procedures.
 Across two benchmarks and two model families, models achieved high predictive performance.
-However, their explanations frequently diverged into multiple distinct modes, even within the same model class and data split.
+However, their explanations frequently diverged into multiple distinct modes whenever the training pipeline admitted more than one near-equivalent solution, even within the same model class and data split.
 The extent and structure of this multiplicity varied by dataset—for example, explanation basins were more evenly balanced on the Breast Cancer dataset and more asymmetric on COMPAS—but in all cases, high predictive accuracy did not imply a single stable mechanism.
@@
 EvoXplain offers a diagnostic view rather than a decision rule.
-It shows that multiplicity exists, shows its form, and specifies the frequency at which different explanatory mechanisms can emerge.
-It does not dictate which local minima should be prioritized; that is a decision to be made from the domain knowledge, validation or regulatory governance.
+It shows that multiplicity exists, shows its form, and specifies the frequency at which different explanatory mechanisms can emerge under admissible retraining.
+It does not dictate which admissible solution should be prioritized; that is a decision to be made from domain knowledge, validation or regulatory governance.
 It is a valuable tool to make multiplicity non-abstract and measurable, and to encourage interpretability practices that account for stability.
@@
 Interpretability can no longer centre on a single trained model.
-The assumed single explanation is viewed as a distribution of explanations, and their stability, diversity, and governance relevance are not assumptions, but empirical matters that require measurement.
+The assumed single explanation should be viewed as a distribution of explanations induced by the training pipeline, and their stability, diversity, and governance relevance are not assumptions, but empirical matters that require measurement.
 EvoXplain pushes the field in this direction by treating explanation stability as a primary diagnostic of model behaviour.
 By making explanation multiplicity observable and quantifiable, EvoXplain provides a foundation for future interpretability, auditing, and governance frameworks that operate at the level of model populations rather than individual fits.
+Appendix A Supplementary Materials and Reproducibility
+All supplementary results, including robustness checks, synthetic controls, and full clustering outputs, are provided in the public code repository.111https://github.com/bensmailchama-boop/EvoXplain/
+The repository contains result folders generated directly by the EvoXplain pipeline, including: (i) per-split clustering summaries, (ii) silhouette and entropy values, (iii) PCA visualisations with basin overlays, and (iv) metadata files specifying experimental configurations.
 Acknowledgements